### v1

# MetaAligner: Conditional Weak-to-Strong Correction for Generalizable Multi-Objective Alignment of Language Models

Abstract
Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses. MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 11 policy models with up to 63$\times$ more parameters, and outperforms previous alignment methods with down to 22.27$\times$ less computational resources. The model also accurately aligns with unseen objectives, marking the first step towards generalizable multi-objective preference alignment. Models and codes will be available at: https://github.com/SteveKGYang/MetaAligner
1 Introduction
The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020), where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model (Ouyang et al., 2022).
Algorithm Paradigm Reward-Free Align. Multi-Objective Align. Policy-Agnostic Align. Generalizability
RLHF (Ouyang et al., 2022) PPO
MORLHF (Li et al., 2020) PPO
MODPO (Zhou et al., 2023) DPO
CPO (Guo et al., 2024) SFT, DPO
RiC (Yang et al., 2024) SFT
Aligner (Ji et al., 2024) SFT
MetaAligner SFT
Table 1: A checklist for key characteristics of previous preference alignment methods and MetaAligner. "Align." denotes "Alignment".
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences (Bai et al., 2022; Rame et al., 2024). These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) (Sener and Koltun, 2018; Li et al., 2020; Rame et al., 2024). Due to its substantial computational cost (Li et al., 2020; Rame et al., 2024) and the unstable nature of the PPO (Schulman et al., 2017; Kumar et al., 2020; Rafailov et al., 2024), other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) (Zhou et al., 2023) or supervised fine-tuning (SFT)-based methods (Yang et al., 2024; Guo et al., 2024), which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023; Chiang et al., 2023; Team et al., 2024). (2) all target models are statically aligned on pre-determined (e.g. Helpfulness, Harmlessness, Honesty (Zhou et al., 2023; Guo et al., 2024)) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access. This is achieved by extending the Aligner (Ji et al., 2024) to multi-objective alignment scenarios, which are not directly solvable by Aligner itself. Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives. The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction. That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering. During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections. As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment. The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning (Kojima et al., 2022). The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. A checklist of MetaAligner’s key advantages over previous methods is presented in Table 1.
Through this work, we make three main contributions:
  • We propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or even accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods with down to 22.27× less computational resources and higher stability.
  • We explore utilizing MetaAligner to exert zero-shot preference alignment for unseen objectives. To the best of our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner performs effective alignment for 3 unseen objectives while maintaining performance on aligned objectives.
  • We examine MetaAligner on the IMHI benchmark (Yang et al., 2023b), which consists of 9 sub-tasks for mental health analysis and 20K testing samples. Experimental results show that MetaAligner trained on merely 7.2K human preference data improves win-rate on multiple objectives across 11 policy models with sizes ranging from 2B to 70B, substantially enhancing responses from policy models with up to 63$\times$ more parameters.
Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF (Li et al., 2020), (B) SFT-based methods (Yang et al., 2024; Guo et al., 2024), and (C) Meta-Objective Aligner. SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
2 Multi-Objective Alignment
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplay such as contradiction (e.g. helpfulness and harmlessness) and dependence (e.g. correctness and informativeness). Multi-objective preference alignment tackles this challenge by optimizing on multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$. $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry (Bradley and Terry, 1952) model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\frac{1}{1+exp(-(r_{i}(q,y_{1})-r_{i}(q,y_{2})))}$ (1)
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards (Sener and Koltun, 2018; Li et al., 2020) with a KL-divergence regularization (see Figure 1 (A)):
$\small\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector. In another paradigm, Yang et al. (2024); Guo et al. (2024) directly build alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q,\mathbf{R}(q,y))\right]$ (3)
where objectives and their corresponding reward values are described with text markers and combined into queries with a static prompting template. Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable.
3 Meta-Objective Aligner
This section introduces the proposed methodology for MetaAligner, which follows a three-stage paradigm: (1) Dynamic multi-objective dataset construction; (2) a three-step model training process; (3) policy-agnostic and generalizable inference for alignment. The paradigm is illustrated in Figure 1 (C).
3.1 Dynamic Multi-Objective Dataset
We propose constructing a dynamic multi-objective dataset to trigger MetaAligner’s ability for reward-free and flexible adjustment of alignment objectives. Specifically, a typical multi-objective preference alignment dataset with $N$ objectives $\mathcal{D}_{n}$ is formalized as $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$, where $R_{i}=[r_{i1},...,r_{iN}]^{\mathrm{T}}$ and $r_{ij}\in\{\succ,\prec,=\}$. We define a text marker for each objective: $<R>:[<r_{1}>,...,<r_{N}>]$, where $<r_{j}>$ denotes a natural language description marking the $i$-th objective. Six examples of such text markers are presented in Figure 1 (C). With a pre-defined prompting template $\mathcal{T}(q,y,\mathcal{O},t)$, we build a preference subset $\mathcal{D}_{p}$ and another equal-preference subset $\mathcal{D}_{e}$ from $\mathcal{D}_{n}$, where $\mathcal{D}_{p}$ includes all contrastive response pairs and $\mathcal{D}_{e}$ includes all equal-labelled response pairs. We utilize the following prompting template for building the datasets:
[$\mathcal{T}(q,y,\mathcal{O},t)$] Edit the following Question-Answer pair to make it {$t$} considering the following objectives $\mathcal{O}$} | Question: {$q$} | Answer: {$y$} | Edit:
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text markers for the target objectives, and $t\in\{equal,better\}$ depends on the building subset. The detailed construction process is described in Algorithm 1.
Algorithm 1 Training data construction
1:Raw datasets: $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$; Objective text markers: $[<r_{1}>,...,<r_{N}>]$; Prompting template: $\mathcal{T}(q,y,\mathcal{O},t)$
2:Preference subset $\mathcal{D}_{p}$; All-equal subset $\mathcal{D}_{e}$.
3:$\mathcal{D}_{p}\leftarrow\varnothing$, $\mathcal{D}_{e}\leftarrow\varnothing$triangleright Initialize the 2 subsets.
4:for $i\in\{1,...,n\}$ do
5:    $\mathcal{O}_{\succ}\leftarrow\varnothing$, $\mathcal{O}_{\prec}\leftarrow\varnothing$, $\mathcal{O}_{=}\leftarrow\varnothing$
6:    for $j\in\{1,...,N\}$ do
7:         if $r_{ij}$ is $\succ$ $\triangleright$ Preferred obj. for $y_{1}$.
8:             $\mathcal{O}_{\succ}\leftarrow\mathcal{O}_{\succ}\cup\{<r_{j}>\}$
9:         else if $r_{ij}$ is $\prec$ $\triangleright$ Preferred obj. for $y_{2}$.
10:             $\mathcal{O}_{\prec}\leftarrow\mathcal{O}_{\prec}\cup\{<r_{j}>\}$
11:         $\triangleright$ Equally preferred objectives.
12:             $\mathcal{O}_{=}\leftarrow\mathcal{O}_{=}\cup\{<r_{j}>\}$
13:         end if
14:    end for
15:    if $\mathcal{O}_{\succ}\neq\varnothing$ $\triangleright$ Build $y_{1}$-preferred data.
16:         $t\leftarrow<better>$
17:         $\mathcal{O}_{\succ}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\succ})$
18:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\succ},t),y_{i1})\}$
19:    end if
20:    if $\mathcal{O}_{\prec}\neq\varnothing$ $\triangleright$ Build $y_{2}$-preferred data.
21:         $t\leftarrow<better>$
22:         $\mathcal{O}_{\prec}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\prec})$
23:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i1},\mathcal{O}_{\prec},t),y_{i2})\}$
24:    end if
25:    if $\mathcal{O}_{=}\neq\varnothing$ $\triangleright$ Build equal data.
26:         $t\leftarrow<equal>$
27:         $\mathcal{O}_{=}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{=})$
28:         $\mathcal{D}_{e}\leftarrow\mathcal{D}_{e}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{=},t),y_{i1})\}$
29:    end if
30:end for
Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping (Yang et al., 2024) process in previous SFT-based multi-objective alignment methods.
3.2 MetaAligner Derivation
Based on the dynamic multi-objective dataset, we achieve multi-objective alignment by introducing MetaAligner, a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the initial outputs as follows:
$\pi^{*}(y|q)=\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\pi_{\phi}(y_{0}|q)$ (4)
where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$. We achieve alignment by directly employing SFT on the target response. We have the standard cross-entropy loss as the training objective:
$\mathop{argmin}\limits_{\theta,\phi}$ $\mathcal{L}(\theta,\phi;\mathcal{D})=-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\pi^{*}(y|q)\right]$ (5)
=-= - $\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-$
$\mathbb{E}_{q\sim\mathcal{D}}\left[log\pi_{\phi}(y_{0}|q)\right]$
We fuse the parameters of the policy model, thus excluding $\phi$ from the weight update process. As only $\theta$ derivatives are calculated during back-propagation, we eliminate the second term in Eqn 5 and simplify the training objective as:
$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as output from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs (Achiam et al., 2023; Team et al., 2023).
3.3 Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a three-step paradigm based on the objective function in Eqn. 6:
Step 1: warm-up.
The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies (He et al., 2016; Ji et al., 2024) to familiarize the model with identity mapping and facilitate convergence. Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
Step 2: equal-preference modelling.
Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, in objectives such as Correctness and Honesty (Yang et al., 2023b; Li et al., 2023b), equal preferences are common and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
Step 3: preference alignment
This stage fine-tunes the MetaAligner on the preference subset $\mathcal{D}_{p}$, which instructs the model to perform conditional weak-to-strong correction on the dis-preferred response and achieve alignment to the preferred response on the specified objectives.
3.4 Generalizable Inference
During inference, MetaAligner achieves alignment following the sampling process as in Eqn. 4, where unaligned outputs, sampled from the target policy model, are used as the input for conditional weak-to-strong correction. With the prompting-based paradigm, the target objectives for MetaAligner also become expandable and generalizable, a key advantage over previous alignment methods (Zhou et al., 2023; Yang et al., 2024; Guo et al., 2024). The generalizability is two-fold: Firstly, users can manipulate the target objectives by adjusting combinations of text markers in the objective set $\mathcal{O}$. For example, in alignment with objectives 1 and 3, we can incorporate the corresponding text markers as follows:
$\mathcal{O}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:<r_{1}>}$ (7)
Secondly, the prompt-based objectives statement enables flexible adjustment of text markers for existing objectives and injections of unseen objectives. Following the last example, we have two unseen alignment objectives $r_{1}^{*}:<r_{1}^{*}>$, $r_{2}^{*}:<r_{2}^{*}>$, and an updated text marker $<r_{1}^{*}>$ for aligned objective 1. We can perform zero-shot alignment on these objectives by adjusting $\mathcal{O}$ as follows:
$\mathcal{O}^{*}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:}{\color[rgb]{1,0,0}<r_{1}^{*}>;r_{1}^{*}:<r_{1}^{*}>;r_{2}^{*}:<r_{2}^{*}>}$ (8)
This simple pattern can theoretically lead to unlimited simultaneous alignment objectives. We expect MetaAligner to make generalizable weak-to-strong corrections under these unseen conditions via its in-context learning ability. This advancement marks a new exploration into generalizable multi-objective preference alignment.
MetaAligner Policy Model CAMS DR Dreaddit IRF loneliness MultiWD SAD SWMH T-SID Overall
1.1B LLaMA2-Chat-7B -3.4% +28.0% +12.67% +45.67% +38.0% +34.67% +23.33% +33.0% +31.33% +27.04%
LLaMA2-Chat-13B -25.0% +27.67% +22.67% +52.33% +32.0% +31.33% +23.33% +23.67% +36.67% +29.7%
LLaMA2-Chat-70B +5.33% +35.0% +19.0% +46.33% +42.0% -3.0% +8.0% +0.33% +3.0% +21.18%
Gemma-instruct-2B +32.0% +4.33% +37.0% +40.0% +34.0% +61.0% +45.0% +52.0% +27.33% +38.77%
Gemma-instruct-7B +16.33% +51.33% +42.67% +44.67% +51.0% +55.33% +49.33% +40.0% +53.67% +44.92%
MentaLLaMA-7B +13.33% +22.67% +23.33% +47.33% +39.67% +39.33% +33.33% +30.67% +41.0% +32.29%
MentaLLaMA-13B +21.34% +31.0% +39.0% +29.67% +42.33% +47.66% +20.67% +34.33% +33.0% +36.03%
MentaLLaMA-33B -5.44% -0.33% -5.33% +30.67% +3.0% +14.0% +24.33% +7.0% +2.33% +6.66%
Vicuna-7B +6.0% -9.33% +24.0% +70.67% +17.67% +33.33% +17.33% +22.0% +42.0% +24.85%
Vicuna-13B +6.0% -9.0% +13.67% +71.67% +20.33% +32.67% +21.0% +2.0% +16.67% +24.97%
Vicuna-33B +29.33% +2.0% -1.0% +41.0% +10.0% +9.67% +1.33% -9.67% +1.33% +9.33%
7B LLaMA2-Chat-7B +7.67% +47.0% +13.0% +33.0% +38.33% +31.0% +17.0% +38.67% +49.33% +30.55%
LLaMA2-Chat-13B -12.34% +51.67% +18.0% +37.67% +43.33% +29.33% +22.33% +18.0% +38.34% +27.37%
LLaMA2-Chat-70B +15.66% +40.53% +9.0% +28.33% +25.34% +6.67% +31.33% +10.33% +10.66% +20.67%
Gemma-instruct-2B +45.0% +8.33% +30.0% +57.0% +39.0% +48.66% +57.0% +52.0% +36.33% +41.11%
Gemma-instruct-7B +32.0% +42.67% +20.67% +45.0% +36.33% +43.0% +46.0% +40.67% +20.7% +38.0%
MentaLLaMA-7B +20.66% +50.67% +19.66% +35.66% +30.0% +27.0% +40.33% +36.34% +41.0% +35.48%
MentaLLaMA-13B +25.0% +45.66% +43.66% +34.34% +36.33% +30.33% +48.0% +42.33% +36.0% +37.97%
MentaLLaMA-33B -5.33% +2.66% -9.66% +2.34% +21.33% -4.67% +20.0% +8.67% +3.33% +4.22%
Vicuna-7B +22.33% +7.67% +4.0% +40.0% +20.0% +14.0% +6.0% +7.67% +15.0% +15.19%
Vicuna-13B +1.0% +48.0% +8.0% +67.0% +33.0% +12.0% +15.0% +23.0% +22.0% +32.33%
Vicuna-33B -2.0% +54.0% +10.0% +62.0% +37.0% +29.0% +10.0% +15.0% +12.0% +25.22%
13B LLaMA2-Chat-7B +27.33% +45.0% +23.33% +62.67% +54.33% +59.34% +52.0% +55.33% +70.33% +49.96%
LLaMA2-Chat-13B +5.34% +47.0% +22.67% +65.33% +56.33% +52.33% +31.33% +32.0% +56.67% +35.55%
LLaMA2-Chat-70B +28.0% +59.0% +23.33% +57.33% +53.0% +7.0% +27.33% +23.0% +23.67% +31.18%
Gemma-instruct-2B +52.66% +19.66% +41.33% +60.67% +50.67% +79.0% +53.67% +63.33% +45% +51.74%
Gemma-instruct-7B +35.33% +52.0% +39.0% +55.34% +50.0% +64.0% +56.33% +49.33% +61.0% +50.36%
MentaLLaMA-7B +40.67% +38.34% +29.33% +65.67% +38.34% +57.0% +46.66% +49.0% +61.67% +47.4%
MentaLLaMA-13B +34.0% +27.0% +39.0% +54.33% +35.33% +49.33% +41.67% +46.67% +45.66% +43.62%
MentaLLaMA-33B +2.67% -8.0% +2.67% +30.67% -8.0% +23.67% +29.33% +28.67% +20.0% +11.2%
Vicuna-7B +23.33% +17.67% +8.0% +63.67% +35.67% +36.33% +12.33% +18.0% +37.0% +28.07%
Vicuna-13B +15.0% -7.33% +2.0% +67.67% +24.33% +40.67% +19.0% +4.0% +17.67% +20.29%
Vicuna-33B -6.0% +54.0% +36.0% +79.0% +38.0% +29.0% +24.0% +21.0% +24.0% +33.22%
Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models. The GPT-4 judge considers 3 objectives: Correctness, Informativeness, and Professionality. The figures show the advantage of aligned outputs over the policy model outputs on win rate. Best values for each MetaAligner model are highlighted in bold.
4 Experiments
4.1 Experimental Settings
Datasets
We train and evaluate the performance of MetaAligner on the IMHI dataset (Yang et al., 2023b). For the training data, we invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives with the following text markers: (1) Correctness: the response should make correct predictions in the corresponding sub-tasks; (2) informativeness: the response should express clear logic and provide consistent supporting evidence; (3) Professionality: the generated responses should provide supportive evidence with high quality and reliability. For evaluation, we perform alignment on the IMHI benchmark, which consists of 9 sub-tasks with 20K query-response pairs.
Models
We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B (Zhang et al., 2024) and LLaMA2-Chat-(7B, 13B) (Touvron et al., 2023) foundation models. We utilize MetaAligner to perform multi-objective alignment on the following policy models: LLaMA2-Chat-(7B,13B,70B) (Touvron et al., 2023), Gemma-instruct-(2B,7B) (Team et al., 2024), MentaLLaMA-(7B, 13B, 33B) (Yang et al., 2023b), and Vicuna-(7B, 13B, 33B) (Chiang et al., 2023). Among the policy models, MentaLLaMA models are instruction-tuned on mental health-related data, and other models are tuned on general-domain instruction data.
Evaluation Metric
We quantify the alignment performance of model outputs by comparing their win rate against the ground-truth response provided by the IMHI benchmark, on the corresponding objectives. Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works (Guo et al., 2024; Sun et al., 2023; Li et al., 2023a), to perform the judgments, where each target response, query, ground-truth response, and objectives are provided via prompt engineering. GPT-4 is required to select the response with higher alignment with the specified objectives in the prompt, or indicate a tied performance.
Figure 2: Objective-wise win-rate comparisons based on different policy models. MetaAligner-7B with different levels of alignment are tested on Correctness, Informativeness, and Professionality, compared with 3 different policy models. The win rates are obtained via the judgments of GPT-4 on the corresponding objectives.
4.2 Overall Performance
We stack MetaAligner on different policy models to perform alignment on all 3 objectives: Correctness, Informativeness, and Professionality. The overall performance on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 2.
According to the results, the MetaAligner models achieve substantial improvement in overall performance on all 11 policy models, with an average of 26.89% advantage on win rates for MetaAligner-1.1B, 28.01% for MetaAligner-7B, and 36.6% for MetaAligner-13B. These results show the general effectiveness of one MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment. MetaAligner also greatly improves performance on each sub-task. For example, MetaAligner-7B outperforms the unaligned outputs by over 25% on 7 sub-tasks. These results indicate that MetaAligner alignment can be effectively adapted to tasks that require different knowledge and response formats.
From the policy model scale perspective, MetaAligner provides successful alignments to models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself. In the extreme case, MetaAligner-1.1B advances the win-rate of LLaMA2-Chat-70B outputs by 21.18%, a policy model with 63$\times$ more parameters. These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow. Besides the general-domain foundation models, MetaAligner also improves the performance by an average of 28.32% on MentaLLaMA models, which are fine-tuned on mental health analysis tasks. These results show that MetaAligner can make reasonable corrections on weak responses while maintaining their expertise from domain-specific policy models.
Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase. These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner. Fortunately, MetaAligner’s capabilities also show scalability. Increasing the size of its base model leads to a higher win-rate advantage on most policy models. Specifically, MetaAligner-7B outperforms or achieves comparable performance to MetaAligner-1.1B on 9 policy models, and MetaAligner-13B surpasses MetaAligner-7B on 10 policy models by an average of 12.58%. These observations motivate further explorations in model scale-performance balance for MetaAligner.
Figure 3: Generalizable alignment on 3 unseen objectives based on LLaMA2-Chat-70B outputs. In the x-axis, "+" denotes the addition of a new objective into the objective set $\mathcal{O}$. "★" denotes the unaligned win rates for LLaMA2-Chat-70B outputs. "-." lines identify win rate fluctuations before alignment on the corresponding objective, and solid lines identify fluctuations after alignment.
4.3 Objective-Wise Evaluation
We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives. We utilize MetaAligner to perform 5 levels of alignments: No alignment, aligning on each objective (Correctness, Informativeness, and Professionality), and full alignment. The objective-wise win rates are shown in Figure 2. We have the following observations:
Objective-wise alignment significantly improves the target objective and benefits the performance on other objectives. For example, aligning on Informativeness with MetaAligner achieves the best performance on Informativeness, and increases win rates on Correctness and Professionality over the unaligned outputs. This tendency holds with all policy models and other objectives. These results further prove the complex interplay within multi-objective alignment, where positive correlations also exist, in addition to contradictions (Guo et al., 2024).
Aligning on all objectives provides balanced performance. According to the results, full alignment displays competitive performance on all 3 objectives and generally outperforms unaligned outputs and aligned outputs from other objectives, even comparable to aligned outputs from the same objective in some cases. The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, enabling the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
MetaAligner performs comprehensive alignment on different sub-tasks. With most policy models, MetaAligner significantly outperforms the unaligned responses on all 9 sub-tasks, regardless of the alignment objectives. These results indicate the general effectiveness of our dynamic-objective method on tasks demanding various expertise (Yang et al., 2023b). However, MetaAligner still struggles to align certain sub-tasks, especially with strong policy models. For example, MetaAligner outputs fail to outperform MentaLLaMA-13B on the CAMS sub-task in all objectives. These results show enhancements in domain knowledge, with techniques such as retrieval-augment generation (Lewis et al., 2020), a promising future direction for improvement.
Algorithm Space Complexity Correctness Informativeness Professionality Overall
MORLHF 22.27$\times$ -6.31% +33.7% +16.31% +15.08%
MODPO 17.27$\times$ +26.31% +30.1% +33.12% +32.9%
SFT 11.82$\times$ +19.22% +24.89% +25.6% +26.62
MetaAligner-1.1B 1$\times$ +26.11% +35.67% +13.11% +29.7%
MetaAligner-7B 6.36$\times$ +28.89% +24.92% +29.07% +27.37%
MetaAligner-13B 11.82$\times$ +31.33% +45.0% +43.66% +35.55%
Table 3: Win rates of previous alignment methods and MetaAligner models based on the LLaMA2-Chat-13B policy model. Space complexity is calculated based on the memory consumption of training MetaAligner-1.1B.
4.4 Generalizable Alignment
In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 3 less relevant objectives that are unseen during its training phases: (1) Completeness (Yang et al., 2023a): the response should cover all relevant aspects of the original query; (2) Honesty (Guo et al., 2024): the response should speak in truth that is consistent with the query and commonsense. (3) Safety (Sun et al., 2023): The response should avoid content that is offensive, discriminatory, or harmful.
2,700 random-selected responses from the IMHI benchmark are re-aligned with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one. Their performance is presented in Figure 3. According to the results, we have the following conclusions:
MetaAligner performs effective zero-shot alignment for unseen objectives. With all MetaAligner models, incorporating an unseen objective significantly improves its corresponding win rate. For example, MetaAligner-7B improves win rate by 40.7% on Completeness, 28.89% on Honesty, and 17.56% on Safety compared to each of these objectives not introduced. These results prove the viability of generalizable alignment with the in-context learning ability of MetaAligner. In addition, alignment performance on supervised objectives (Correctness, Informativeness, Professionality) generally surpasses unseen objectives on all tested models, showing that MetaAligner also learns objective-specific features during training.
Performance on aligned objectives is maintained with more unseen alignment objectives. As each objective is aligned, its win rate surges, then stabilizes as long as the objective is included in $\mathcal{O}$. These results prove MetaAligner to perform reliable alignment with the expansion of objectives. However, aligning on new objectives can affect other unaligned objectives. For example, aligning on Correctness with MetaAligner-1.1B benefits its win rate by 11.22%, but harms performance on 3 other unaligned objectives. These problems are due to the limited ability of smaller MetaAligner models and the controversial dependence of certain multi-objective alignments (Guo et al., 2024), where we observe an alleviation as the model size increases.
MetaAligner’s generalizability also shows scalability. Performance on the 3 unseen objectives increases with the scale-up of MetaAligner model size, where MetaAligner-1.1B improves 12.7% on average, MetaAligner-7B improves 29.05%, and MetaAligner-13B improves 36.69%. This scalability is attributed to the growing in-context learning ability of larger base models, which motivates further explorations into the correlation between generalizable alignment and model scales.
4.5 MetaAligner vs. MORLHF/MODPO/SFT
We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on 2700 random-selected responses from the IMHI benchmark. We implement the linear scalarization method for MORLHF. We follow previous settings of realizing MODPO (Guo et al., 2024) by pre-pending the objective markers. Due to the closed-source nature of current SFT-based methods, we only replicate stage 1 where details are clear, and utilize the preference-to-reward mapping method proposed in Yang et al. (2024). We select LLaMA2-Chat-13B as the policy model, and the results are presented in Table 3.
According to the results, MetaAligner-13B achieves over 30% of win rates on all objectives, and generally outperforms all previous methods by 10.68% on average. This advantage is achieved with equal or less space complexity than previous algorithms, showing the high efficiency of the proposed method. As the base model size further reduces, MetaAligner-1.1B and MetaAligner-7B show lower, but still competitive performance to the 13B model. Impressively, MetaAligner-1.1B outperforms both MORLHF and SFT-based methods by 8.85% on average, and approaches MODPO with merely a 3.2% gap, but with 11.82$\times$-22.27$\times$ less computational resources. This huge advantage indicates a high prospect for applying MetaAligner in low-resource scenarios. For previous methods, MODPO and SFT can provide stable alignment for multiple objectives, but with high space complexity. MORLHF bears the highest resource consumption, but achieves the lowest 15.08% overall win rate. Its training process is also unstable, including a 6.31% decrease in the Correctness objective.
5 Related Work
5.1 Large Language Models
Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language Achiam et al. (2023); Team et al. (2023); Touvron et al. (2023); Luo et al. (2023). Notable examples include commercially available LLMs like ChatGPT, GPT-4 Achiam et al. (2023), and Gemini Team et al. (2023), as well as leading-edge open-source models like LLaMA2 Touvron et al. (2023) and Mistral Jiang et al. (2023). Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information Zhang et al. (2023); Xie et al. (2023), flattery, and deception, raising concerns about their potential negative impacts on society and associated risks Bommasani et al. (2021). To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning Ji et al. (2023).
5.2 Preference Alignment
Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty. These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022). Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences. This limitation has motivated the exploration of multi-objective RLHF (MORLHF) Sener and Koltun (2018); Li et al. (2020); Rame et al. (2024), by linear scalarizations of multiple rewards Sener and Koltun (2018); Li et al. (2020) and interpolations of LLM weights trained from diverse reward models Rame et al. (2024). However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) Zhou et al. (2023) without the RL paradigm. Zhou et al. (2023) proposed multi-objective direct preference optimization (MODPO), which extended DPO algorithm (Rafailov et al., 2024) to combine multiple objectives with specific weightings. Another paradigm is the SFT-based methods Yang et al. (2024); Guo et al. (2024), which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO. These methods also facilitated objective-wise controllable generation during inference.
6 Discussion
Conclusion.
This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment. Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives. Thorough investigations on LLMs ranging from 2B to 70B proved MetaAligner’s overall effectiveness in multi-objective alignment for various policy models and objective-wise alignment. and Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align more objectives.
Limitation.
Though achieving outstanding performance, MetaAligner still bears limitations. Firstly, similar to the vanilla Aligner module Ji et al. (2024), stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference, which affects actual model deployment, especially for scenarios such as local deployment on mobile devices. Secondly, MetaAligner still provides limited improvement for large-scale domain-specific policy models, such as MentaLLaMA-33B, in domain knowledge-intense tasks such as mental health analysis. These results reflect the importance of expertise knowledge for multi-objective alignment, where MetaAligner still requires further enhancements. Thirdly, due to limited resources, we only tested the generalizability of MetaAligner on 3 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives.
Future Work.
Firstly, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation (Lewis et al., 2020). Secondly, we will dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance. We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 6 objectives). This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.

### v2

@@
-MetaAligner: Conditional Weak-to-Strong Correction for Generalizable Multi-Objective Alignment of Language Models
+MetaAligner: Towards Generalizable Multi-Objective Alignment of Language Models
 Abstract
@@
 MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning.
-Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 11 policy models with up to 63$\times$ more parameters, and outperforms previous alignment methods with down to 22.27$\times$ less computational resources.
-The model also accurately aligns with unseen objectives, marking the first step towards generalizable multi-objective preference alignment.
-Models and codes will be available at: https://github.com/SteveKGYang/MetaAligner
+Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and outperforms previous alignment methods with down to 15.71$\times$ less GPU training hours.
+The model also effectively aligns unseen objectives, marking the first step towards generalizable multi-objective preference alignment.
+This project is open-sourced at: https://github.com/SteveKGYang/MetaAligner
 1 Introduction
 The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values.
-At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020), where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model (Ouyang et al., 2022).
+At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [39, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model [23].
+However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences [4, 25].
+These drawbacks motivate further exploration into multi-objective alignment algorithms.
+Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) [27, 20, 25].
+Due to its substantial computational cost [20, 25] and the unstable nature of the PPO [26, 17, 24], other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [38, 11] or supervised fine-tuning (SFT)-based methods [35, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
+The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model.
+This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g.
+"Helpful", "Harmless", "Honest" [38, 11]) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives.
+This ignorance leads to poor generalizability of existing multi-objective alignment methods.
 Algorithm Paradigm Reward-Free Align. Multi-Objective Align. Policy-Agnostic Align. Generalizability
-RLHF (Ouyang et al., 2022) PPO
-MORLHF (Li et al., 2020) PPO
-MODPO (Zhou et al., 2023) DPO
-CPO (Guo et al., 2024) SFT, DPO
-RiC (Yang et al., 2024) SFT
-Aligner (Ji et al., 2024) SFT
+RLHF [23] PPO
+MORLHF [20] PPO
+MODPO [11, 38] SFT, DPO
+RiC [35] SFT
+Aligner [13] SFT
 MetaAligner SFT
 Table 1: A checklist for key characteristics of previous preference alignment methods and MetaAligner. "Align." denotes "Alignment".
-However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences (Bai et al., 2022; Rame et al., 2024).
-These drawbacks motivate further exploration into multi-objective alignment algorithms.
-Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) (Sener and Koltun, 2018; Li et al., 2020; Rame et al., 2024).
-Due to its substantial computational cost (Li et al., 2020; Rame et al., 2024) and the unstable nature of the PPO (Schulman et al., 2017; Kumar et al., 2020; Rafailov et al., 2024), other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) (Zhou et al., 2023) or supervised fine-tuning (SFT)-based methods (Yang et al., 2024; Guo et al., 2024), which customized prompting strategies to incorporate multiple reward values into queries explicitly.
-The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model.
-This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023; Chiang et al., 2023; Team et al., 2024).
-(2) all target models are statically aligned on pre-determined (e.g.
-Helpfulness, Harmlessness, Honesty (Zhou et al., 2023; Guo et al., 2024)) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives.
-This ignorance leads to poor generalizability of existing multi-objective alignment methods.
 In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment.
-MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access.
-This is achieved by extending the Aligner (Ji et al., 2024) to multi-objective alignment scenarios, which are not directly solvable by Aligner itself.
+MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access [13].
 Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives.
@@
 As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment.
-The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning (Kojima et al., 2022).
+The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning [16].
 The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives.
-A checklist of MetaAligner’s key advantages over previous methods is presented in Table 1.
-Through this work, we make three main contributions:
-  • We propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment.
-It performs multi-objective alignment efficiently, without tuning the policy models or even accessing their parameters.
-Experimental results show that MetaAligner outperforms previous alignment methods with down to 22.27× less computational resources and higher stability.
-  • We explore utilizing MetaAligner to exert zero-shot preference alignment for unseen objectives.
-To the best of our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment.
-Experimental results show that MetaAligner performs effective alignment for 3 unseen objectives while maintaining performance on aligned objectives.
-  • We examine MetaAligner on the IMHI benchmark (Yang et al., 2023b), which consists of 9 sub-tasks for mental health analysis and 20K testing samples.
-Experimental results show that MetaAligner trained on merely 7.2K human preference data improves win-rate on multiple objectives across 11 policy models with sizes ranging from 2B to 70B, substantially enhancing responses from policy models with up to 63$\times$ more parameters.
-Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF (Li et al., 2020), (B) SFT-based methods (Yang et al., 2024; Guo et al., 2024), and (C) Meta-Objective Aligner.
+A checklist of MetaAligner’s advantages over previous methods is presented in Table 1.
+Through this work, we make three main contributions: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment.
+It performs multi-objective alignment efficiently, without tuning the policy models or accessing their parameters.
+Experimental results show that MetaAligner outperforms previous alignment methods with down to 15.71× less GPU training hours and higher stability; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives.
+To our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment.
+Experimental results show that MetaAligner can simultaneously perform effective alignment for 6 unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on 3 preference alignment datasets.
+Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5 [22] and Claude-3 [3].
+Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF [20], (B) SFT-based methods [35, 11], and (C) Meta-Objective Aligner.
 SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
-2 Multi-Objective Alignment
-In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplay such as contradiction (e.g.
-helpfulness and harmlessness) and dependence (e.g.
-correctness and informativeness).
-Multi-objective preference alignment tackles this challenge by optimizing on multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$.
-$r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry (Bradley and Terry, 1952) model of preferences.
+2 Multi-Objective Alignment of Language Models
+In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplays such as contradiction (e.g.
+"Helpful" and "Harmless") and dependence (e.g.
+"Correct" and "Informative").
+Multi-objective preference alignment tackles this challenge by optimizing multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$.
+$r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences.
 Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
@@
 where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$.
-MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards (Sener and Koltun, 2018; Li et al., 2020) with a KL-divergence regularization (see Figure 1 (A)):
-$\small\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
+MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization (see Figure 1 (A)):
+$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
 where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector.
-In another paradigm, Yang et al.
-(2024); Guo et al.
-(2024) directly build alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
+Another paradigm [35, 11] directly built alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
 $\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q,\mathbf{R}(q,y))\right]$ (3)
@@
 3 Meta-Objective Aligner
-This section introduces the proposed methodology for MetaAligner, which follows a three-stage paradigm: (1) Dynamic multi-objective dataset construction; (2) a three-step model training process; (3) policy-agnostic and generalizable inference for alignment.
+This section introduces the proposed methodology for MetaAligner, which follows a 3-stage paradigm: (1) dynamic multi-objective dataset construction; (2) a 3-step model training process; (3) generalizable inference for alignment.
 The paradigm is illustrated in Figure 1 (C).
 3.1 Dynamic Multi-Objective Dataset
-We propose constructing a dynamic multi-objective dataset to trigger MetaAligner’s ability for reward-free and flexible adjustment of alignment objectives.
-Specifically, a typical multi-objective preference alignment dataset with $N$ objectives $\mathcal{D}_{n}$ is formalized as $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$, where $R_{i}=[r_{i1},...,r_{iN}]^{\mathrm{T}}$ and $r_{ij}\in\{\succ,\prec,=\}$.
-We define a text marker for each objective: $<R>:[<r_{1}>,...,<r_{N}>]$, where $<r_{j}>$ denotes a natural language description marking the $i$-th objective.
-Six examples of such text markers are presented in Figure 1 (C).
-With a pre-defined prompting template $\mathcal{T}(q,y,\mathcal{O},t)$, we build a preference subset $\mathcal{D}_{p}$ and another equal-preference subset $\mathcal{D}_{e}$ from $\mathcal{D}_{n}$, where $\mathcal{D}_{p}$ includes all contrastive response pairs and $\mathcal{D}_{e}$ includes all equal-labelled response pairs.
-We utilize the following prompting template for building the datasets:
+[…]
 [$\mathcal{T}(q,y,\mathcal{O},t)$] Edit the following Question-Answer pair to make it {$t$} considering the following objectives $\mathcal{O}$} | Question: {$q$} | Answer: {$y$} | Edit:
-where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text markers for the target objectives, and $t\in\{equal,better\}$ depends on the building subset.
+where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the building subset.
 The detailed construction process is described in Algorithm 1.
-Algorithm 1 Training data construction
-1:Raw datasets: $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$; Objective text markers: $[<r_{1}>,...,<r_{N}>]$; Prompting template: $\mathcal{T}(q,y,\mathcal{O},t)$
-2:Preference subset $\mathcal{D}_{p}$; All-equal subset $\mathcal{D}_{e}$.
-3:$\mathcal{D}_{p}\leftarrow\varnothing$, $\mathcal{D}_{e}\leftarrow\varnothing$triangleright Initialize the 2 subsets.
-4:for $i\in\{1,...,n\}$ do
-5:    $\mathcal{O}_{\succ}\leftarrow\varnothing$, $\mathcal{O}_{\prec}\leftarrow\varnothing$, $\mathcal{O}_{=}\leftarrow\varnothing$
-6:    for $j\in\{1,...,N\}$ do
-7:         if $r_{ij}$ is $\succ$ $\triangleright$ Preferred obj. for $y_{1}$.
-8:             $\mathcal{O}_{\succ}\leftarrow\mathcal{O}_{\succ}\cup\{<r_{j}>\}$
-9:         else if $r_{ij}$ is $\prec$ $\triangleright$ Preferred obj. for $y_{2}$.
-10:             $\mathcal{O}_{\prec}\leftarrow\mathcal{O}_{\prec}\cup\{<r_{j}>\}$
-11:         $\triangleright$ Equally preferred objectives.
-12:             $\mathcal{O}_{=}\leftarrow\mathcal{O}_{=}\cup\{<r_{j}>\}$
-13:         end if
-14:    end for
-15:    if $\mathcal{O}_{\succ}\neq\varnothing$ $\triangleright$ Build $y_{1}$-preferred data.
-16:         $t\leftarrow<better>$
-17:         $\mathcal{O}_{\succ}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\succ})$
-18:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\succ},t),y_{i1})\}$
-19:    end if
-20:    if $\mathcal{O}_{\prec}\neq\varnothing$ $\triangleright$ Build $y_{2}$-preferred data.
-21:         $t\leftarrow<better>$
-22:         $\mathcal{O}_{\prec}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\prec})$
-23:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i1},\mathcal{O}_{\prec},t),y_{i2})\}$
-24:    end if
-25:    if $\mathcal{O}_{=}\neq\varnothing$ $\triangleright$ Build equal data.
-26:         $t\leftarrow<equal>$
-27:         $\mathcal{O}_{=}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{=})$
-28:         $\mathcal{D}_{e}\leftarrow\mathcal{D}_{e}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{=},t),y_{i1})\}$
-29:    end if
-30:end for
-Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping (Yang et al., 2024) process in previous SFT-based multi-objective alignment methods.
+A full list of the utilized prompting templates is presented in Appendix B.
+Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping [35] process in previous SFT-based multi-objective alignment methods.
 3.2 MetaAligner Derivation
@@
 $\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
-The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as output from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs (Achiam et al., 2023; Team et al., 2023).
+The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as samples from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs [2, 22, 3].
 3.3 Model Training
 In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process.
-We propose a three-step paradigm based on the objective function in Eqn.
+We propose a 3-step paradigm based on the objective function in Eqn.
 6:
 Step 1: warm-up.
-The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies (He et al., 2016; Ji et al., 2024) to familiarize the model with identity mapping and facilitate convergence.
+The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies [12, 13] to familiarize the model with identity mapping and facilitate convergence.
 Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
-Step 2: equal-preference modelling.
+Step 2: equal-preference modeling.
 Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs.
-However, in objectives such as Correctness and Honesty (Yang et al., 2023b; Li et al., 2023b), equal preferences are common and enclose useful information such as the principle components of preference modeling regarding each objective.
+However, equal preferences are common in many scenarios [34, 8] and enclose useful information such as the principle components of preference modeling regarding each objective.
 Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
@@
 3.4 Generalizable Inference
-During inference, MetaAligner achieves alignment following the sampling process as in Eqn.
-4, where unaligned outputs, sampled from the target policy model, are used as the input for conditional weak-to-strong correction.
-With the prompting-based paradigm, the target objectives for MetaAligner also become expandable and generalizable, a key advantage over previous alignment methods (Zhou et al., 2023; Yang et al., 2024; Guo et al., 2024).
-The generalizability is two-fold: Firstly, users can manipulate the target objectives by adjusting combinations of text markers in the objective set $\mathcal{O}$.
-For example, in alignment with objectives 1 and 3, we can incorporate the corresponding text markers as follows:
-$\mathcal{O}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:<r_{1}>}$ (7)
-Secondly, the prompt-based objectives statement enables flexible adjustment of text markers for existing objectives and injections of unseen objectives.
-Following the last example, we have two unseen alignment objectives $r_{1}^{*}:<r_{1}^{*}>$, $r_{2}^{*}:<r_{2}^{*}>$, and an updated text marker $<r_{1}^{*}>$ for aligned objective 1.
-We can perform zero-shot alignment on these objectives by adjusting $\mathcal{O}$ as follows:
-$\mathcal{O}^{*}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:}{\color[rgb]{1,0,0}<r_{1}^{*}>;r_{1}^{*}:<r_{1}^{*}>;r_{2}^{*}:<r_{2}^{*}>}$ (8)
+[…]
+[…]
+[…]
+[…]
 This simple pattern can theoretically lead to unlimited simultaneous alignment objectives.
@@
 This advancement marks a new exploration into generalizable multi-objective preference alignment.
+4 Experiments
+4.1 Experimental Settings
+Datasets.
+We train and evaluate on 3 alignment datasets: (1) HH-RLHF [4]: a large-scale dataset with 160K prompts and corresponding response pairs.
+We follow previous works [35] and utilize 3 open-sourced reward models on "Harmless", "Helpful", and "Humor" to assess and rank the responses in their corresponding objectives; (2) UltraFeedback [8]: a multi-aspect alignment dataset with 64K prompts with preferences obtained from GPT-4.
+We select "Instruction following", "Honest", "Truthful", and "Helpful" objectives during training; (3) IMHI: we evaluate MetaAligner on domain knowledge-intense scenarios by creating another alignment dataset based on the IMHI dataset [34] targeting interpretable mental health analysis, which consists of 9 mental health analysis sub-tasks.
+We invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives: "Correct", "Informative", and "Professional".
+Models.
+We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B [36] and LLaMA2-(7B, 13B) [31] foundation models.
+We utilize MetaAligner to perform multi-objective alignment on the following open-source policy models: LLaMA2-Chat-(7B,13B,70B) [31], Gemma-instruct-(2B,7B) [30], and Vicuna-(7B, 13B, 33B) [7].
+We also align 2 advanced close-sourced foundation models: GPT-3.5 [22] and Claude-3-Sonnet [3], where model parameters are inaccessible.
+Evaluation Metric.
+On each objective, we quantify the alignment performance of model outputs by comparing their win rates against the ground-truth response provided by the benchmark datasets.
+Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works [11, 29, 19], to perform the judgments.
+Each target response, ground-truth response, query, and evaluated objectives are provided via prompt engineering.
+GPT-4 is required to compare and select the response with higher alignment to the specified objectives or indicate a tied performance.
+More details about the training process, model cards, dataset statistics, IMHI dataset annotation, and evaluation settings are presented in Appendix C.
+HH-RLHF UltraFeedback IMHI
+MetaAligner Policy Model Harmless Helpful Humor IF Honest Truthful Helpful Correct Informative Professional
+1.1B LLaMA2-Chat-7B +10.0% +20.0% +14.75% +11.0% +15.0% +14.33% +9.0% +18.33% +20.55% +31.67%
+LLaMA2-Chat-13B +10.75% +9.08% +13.25% +8.66% +15.34% +16.33% +7.67% +11.11% +8.33% +25.0%
+LLaMA2-Chat-70B +6.58% +7.42% +22.58% +6.0% +12.67% +17.33% +16.33% +8.33% +14.23% +17.23%
+Gemma-instruct-2B +8.5% +12.25% +12.33% +14.67% +14.67% +13.0% +5.33% 15.55% +35.55% +37.23%
+Gemma-instruct-7B +4.0% +7.75% +23.17% +9.0% +10.0% +4.67% +14.0% +18.9% +31.12% +36.11%
+Vicuna-7B +11.5% +10.83% +20.33% +11.33% +13.33% +12.33% +7.0% +10.0% +7.22% +6.33%
+Vicuna-13B +7.42% +13.0% +19.17% +11.66% +14.34% +15.33% +10.0% +12.22% +7.78% +3.34%
+Vicuna-33B +8.5% +2.59% +23.83% +8.0% +11.67% +6.33% +6.67% +8.34% +4.44% +6.12%
+GPT-3.5 +1.42% +7.5% +17.84% +5.0% +5.0% +3.66% +1.0% +9.67% +1.33% +9.33%
+Claude-3-Sonnet -3.83% +1.58% +13.17% +4.67% +2.67% +2.67% +3.0% +7.0% +2.33% +6.66%
+7B LLaMA2-Chat-7B +25.0% +27.0% +20.75% +34.66% +36.0% +37.0% +28.0% +21.67% +32.22% +43.89%
+LLaMA2-Chat-13B +28.75% +20.58% +18.25% 34.0% +37.34% +37.66% +23.3% +25.56% +30.0% +33.89%
+LLaMA2-Chat-70B +16.58% +14.42% +29.08% +31.0% +27.0% +31.33% +17.0% +20.56% +17.23% +21.67%
+Gemma-instruct-2B +20.0% +18.75% +17.83% +41.33% +40.67% +42.33% +31.33% +25.0% +50.55% +51.67%
+Gemma-instruct-7B +11.0% +23.25% +26.67% +33.67% +35.34% +31.0% +29.0% +35.01% +52.23% +56.11%
+Vicuna-7B +19.5% +18.83% +27.33% +38.0% +39.0% +37.0% +32.33% +23.33% +22.78% +23.33%
+Vicuna-13B +14.92% +21.0% +30.67% +34.66% +40.0% +39.67% +36.34% +25.55% +20.0% +15.01%
+Vicuna-33B +28.0% +17.09% +30.83% +30.0% +37.34% +32.33% +29.33% +11.11% +16.11% +8.34%
+GPT-3.5 +15.92% +21.5% +22.84% +29.99% +30.34% +28.0% +14.34% +18.67% +16.33% +14.22%
+Claude-3-Sonnet +19.17% +19.08% +26.17% +22.33% +21.0% +21.67% +19.0% +11.33% +19.33% +11.33%
+13B LLaMA2-Chat-7B +21.0% +30.5% +23.75% +51.83% +47.5% +45.33% +38.67% +28.33% +38.33% +50.56%
+LLaMA2-Chat-13B +22.75% +16.08% +12.25% +46.33% +48.67% +46.83% +41.17% +30.56% +37.22% +40.56%
+LLaMA2-Chat-70B +16.58% +11.92% +23.08% +44.33% +35.0% +45.5% +24.0% +31.67% +30.56% +36.12%
+Gemma-instruct-2B +27.5% +20.75% +24.33% +55.0% +44.67% +51.33% +36.83% +35.55% +63.33% +65.0%
+Gemma-instruct-7B +14.0% +27.75% +25.17% +42.0% +40.17% +35.17% +31.17% +34.45% +50.0% +49.44%
+Vicuna-7B +18.0% +12.83% +30.33% +41.5% +39.83% +44.33% +37.5% +24.44% +23.33% +21.11%
+Vicuna-13B +17.92% +29.0% +32.67% +47.33% +49.17% +47.0% +40.67% +28.33% +23.34% +18.9%
+Vicuna-33B +31.5% +17.09% +27.83% +50.5% +53.17% +45.83% +38.5% +23.89% +23.89% +14.45%
+GPT-3.5 +19.92% +25.0% +23.84% +40.33% +40.17% +36.83% +23.67% +26.67% +25.66% +33.62%
+Claude-3-Sonnet +27.17% +23.58% +27.17% +38.5% +39.5% +37.67% +29.83% +28.67% +20.0% +11.2%
+Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on 3 datasets over different policy models.
+The responses are simultaneously aligned on all objectives, then tested on each objective.
+"IF" denotes the "Instruction following" objective.
+"+" shows the advantage of aligned outputs over the unaligned outputs on win rates against the golden responses.
+4.2 Overall Performance
+MetaAligner-(1.1B, 7B, 13B) performance on 3 alignment datasets are shown in Table 2.
+According to the results, the MetaAligner models achieve substantial improvement for most objectives and policy models.
+For example, on UltraFeedback, there is an average of 11.47% advantage for MetaAligner-1.1B on "Honest", 34.39% for MetaAligner-7B, and 43.79% for MetaAligner-13B.
+These results show the general effectiveness of MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment.
+On the mental health analysis benchmark IMHI, MetaAligner models also show remarkable win rates on all objectives, proving their effectiveness in performing multi-objective alignment in domain knowledge-intense scenarios.
+We further evaluate MetaAligner on each IMHI sub-task and the results are shown in Appendix D.
+From the policy model scale perspective, MetaAligner provides successful alignments to open-sourced models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself.
+In the extreme case, MetaAligner-1.1B advances the win rates of LLaMA2-Chat-70B outputs, a policy model with 63$\times$ more parameters, by an average of 12.19% on HH-RLHF, 13.08% on UltraFedback, and 13.26% on IMHI.
+These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow.
+MetaAligner also significantly improves performance on API-based models: GPT-3.5 and Claude-3-Sonnet.
+These results prove its potential for application in close-sourced scenarios and effective multi-objective alignment of state-of-the-art policy models.
+Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase.
+These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner.
+Fortunately, MetaAligner’s capabilities also show scalability.
+Increasing the size of its base model leads to a higher win-rate advantage on most policy models.
+For example, on UltraFeedback, MetaAligner-7B outperforms MetaAligner-1.1B on all 10 policy models, and MetaAligner-13B further surpasses MetaAligner-7B by an average of 12.58%.
+These observations motivate further explorations in model scale-performance balance for MetaAligner.
+4.3 Evaluations of Objective-Wise Alignment
+Figure 2: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
+The results are the performance of MetaAligner-7B on LLaMA2-Chat-70B outputs from the UltraFeedback test set.
+We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives.
+We utilize MetaAligner to perform 6 levels of alignments: unaligned, aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and full alignment.
+We leverage GPT-4 to score the responses ranging from 0 to 10.
+Experimental details and more results are shown in Appendix E.
+The results are shown in Figure 2.
+We have the following observations:
+Objective-wise alignment improves performance on the target objective and benefits the performance on other objectives.
+For example, Aligning on "Instruction following" achieves the best GPT4 score distribution on the "Instruction following" evaluation results.
+It also significantly increases GPT4 scores on "Helpful", "Honest", and "Truthful" over the unaligned responses.
+This tendency holds with other policy models and alignment objectives.
+These results further prove the complex interplay within multi-objective alignment, where positive correlations and contradictions [11] co-exist.
+Full alignment on all objectives provides balanced performance.
+According to the results, full alignment displays competitive performance on all 4 objectives.
+Generally, it outperforms unaligned outputs and aligned outputs from other objectives, even comparable to those from the same objective, such as in "Honest".
+The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, training the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
+Figure 3: Zero-shot alignment on 6 unseen objectives based on LLaMA2-Chat-70B outputs of UltraFeedback.
+In the x-axis, "Aligned Obj." denotes the 4 supervised objectives ($\diamond$" markers), and "+" denotes further addition of an unseen objective ($\circ$" markers).
+$\star$" denotes the unaligned win rates for the unseen objectives, "-." lines identify win rate fluctuations before alignment, and solid lines identify fluctuations after alignment.
+4.4 Generalizable Alignment
+In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 6 less relevant objectives that are unseen during its training phases: Specific, Factual, Readable, Fair, Repeat, and Length [10].
+More details about these objectives are in Appendix B.
+We randomly select 2,700 queries from the UltraFeedback dataset and re-align the LLaMA2-Chat-70B outputs with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one, with 10 aligned objectives in total.
+Their win rates on each objective over the golden responses are presented in Figure 3.
+According to the results, we have the following conclusions:
+MetaAligner performs effective zero-shot alignment for unseen objectives.
+With most MetaAligner models, incorporating an unseen objective into the objective set significantly improves its corresponding win rate.
+For example, MetaAligner-7B improves by 25.17% on "Specific", 14.5% on "Factual", and 17.5% on "Readable" compared to each of these objectives unaligned.
+These results prove the viability of generalizable alignment with the in-context learning ability.
+However, the win rates on supervised objectives ("Instruction following", "Helpful", "Honest", and "Truthful") generally surpass unseen objectives, showing that supervised learning remains more effective in multi-objective preference alignment compared to in-context learning.
+Performance on aligned objectives is maintained with more unseen alignment objectives.
+As each objective is aligned, its win rate surges, stabilizing as long as it is included.
+On simultaneously aligning 10 objectives, MetaAligner-7B outperforms LLaMA2-Chat-70B outputs by an average of 14.25% on unseen objectives.
+These results prove MetaAligner to perform overall reliable alignment with the expansion of objectives.
+However, aligning on new objectives can affect other objectives.
+For example, aligning on "Fair" with MetaAligner-(7B, 13B) benefits its win rates, but harms performance on objectives such as "Readable" and "Factual".
+The decreases are due to the "alignment tax" [11], where enhancements in one objective can diminish performance in other objectives due to their controversial nature.
+MetaAligner’s generalizability shows scalability.
+Performance on the 6 unseen objectives increases with the scale-up of MetaAligner model size.
+MetaAligner-1.1B provides limited alignment on most unseen objectives, but MetaAligner-7B extends the win rates to an average of 48.5%, and MetaAligner-13B further reaches 61.25%.
+MetaAligner-13B also more effectively aligns objectives such as "Length", where smaller models perform badly.
+This scalability is attributed to larger foundation models’ growing in-context learning ability, which enables accurate interpretations of the objective descriptions and instructions.
+These observations motivate further explorations into the correlation between generalizable alignment and model scales in future work.
+HH-RLHF UltraFeedback
+Algorithm GPU Hours Harmless Helpful Humour Avg. IF Honest Truthful Helpful Avg.
+MORLHF 1892.3 62.83% 51.2% 77.5% 63.84% 32.18% 33.7% 26.1% 33.7% 31.42%
+MODPO 405.9 65.0% 64.0% 78.0% 69.0% 30.82% 43.4% 37.19% 25.0% 34.1%
+SFT 247.34 66.5% 75.0% 76.5% 72.67% 27.0% 36.5% 26.0% 36.5% 31.5%
+MetaAligner-1.1B 120.48 62.5% 75.0% 77.0% 71.5% 27.67% 27.0% 33.0% 25.33% 28.25%
+MetaAligner-7B 242.68 77.5% 82.0% 83.0% 80.83% 51.33% 48.0% 55.67% 44.33% 49.83%
+MetaAligner-13B 403.44 73.5% 85.5% 86.0% 81.67% 68.5% 59.5% 64.0% 55.0% 61.75%
+Table 3: Comparisons of win rates between alignment methods based on the LLaMA2-Chat-7B policy model. "GPU hours" records the summed GPU running time on all datasets.
+4.5 MetaAligner vs. MORLHF/MODPO/SFT-based Method
+We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on the HH-RLHF and UltraFeedback datasets, where explicit rewards are available.
+We implement the linear scalarization method for MORLHF, the CDPO [11] realization of MODPO, and RiC [35] realization of the SFT-based method.
+We select LLaMA2-Chat-7B as the policy model, and the results are presented in Table 3.
+Appendix F presents details about the baseline model implementations and GPU hours calculations.
+According to the results, MetaAligner-13B significantly outperforms all other methods with an average of 81.67% win rate on HH-RLHF and 61.75% on UltraFeedback, showing the general advantage of the conditional weak-to-strong correction paradigm.
+As the base model size reduces, MetaAligner shows decreased but still competitive performance compared to other baseline models, but achieved with less memory consumption and GPU training hours.
+Impressively, the MetaAligner-1.1B model achieves comparable average performance to MORLHF, MODPO, and SFT-based methods on both datasets, but with 2.05$\times$-15.67$\times$ less GPU training hours and 6.36$\times$ smaller size than the LLaMA2-Chat-7B policy model.
+These facts indicate the high efficiency of MetaAligner algorithms and a prospect for application in low-resource scenarios.
+Compared to previous methods, MetaAligner models can also achieve balanced performances in objective-wise evaluations, including contradictory objectives such as "Harmless" and "Helpful", without requiring explicit hyper-parameter tuning for achieving Pareto optimal solutions [11, 35, 20].
+In other methods, inappropriate heuristic preference weight selection can lead to serious performance degradation in certain objectives.
+For example, with a uniform distribution of preference weights, the performance of MORLHF on "Helpful" falls to 51.2%, a huge gap to other methods.
+5 Related Work
+5.1 Large Language Models
+Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language [2, 31, 21].
+Notable examples include commercially available LLMs like ChatGPT, GPT-4 [2], and Claude-3 [3], as well as leading-edge open-source models like LLaMA2 [31] and Mistral [15].
+Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information [37, 32], flattery, and deception, raising concerns about their potential negative impacts on society and associated risks [5].
+To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning [14].
+5.2 Human Preference Alignment
+Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty.
+These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) [23].
+Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences.
+This limitation has motivated the exploration of multi-objective RLHF (MORLHF) [27, 20, 25], by linear scalarizations of multiple rewards [27, 20] and interpolations of LLM weights trained from diverse reward models [25].
+However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) [38] without the RL paradigm.
+Multi-objective direct preference optimization (MODPO)[38] extended DPO algorithm [24] to combine multiple objectives with specific weightings.
+Another paradigm is the SFT-based methods [35, 11], which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO.
+These methods also facilitated objective-wise controllable generation during inference.
+6 Discussions
+Conclusion.
+This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment.
+Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives.
+Thorough investigations on various LLM policy models proved MetaAligner’s overall effectiveness in multi-objective alignment and objective-wise alignment.
+Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align multiple objectives.
+Limitation and Future Work.
+Firstly, stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference [13], which affects model deployment, especially for scenarios such as local deployment on mobile devices.
+Secondly, due to limited resources, we only tested the generalizability of MetaAligner on 6 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives.
+In future work, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation [18].
+We will also dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance.
+We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 10 objectives).
+This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
+Appendix A Ethics and Impacts
+A.1 Licenses
+We leveraged 3 publicly available datasets to build our dynamic multi-objective datasets: HH-RLHF, UltraFeedback, and IMHI.
+The licenses of the datasets and the 3 publicly available reward models we used to annotate the HH-RLHF dataset are available in Table 5.
+In Sec.
+4.5, we implement the reward assignment scripts for HH-RLHF and MORLHF based on the released codes of Yang et al.
+[35], which is available at Github.
+The MODPO and MORLHF codes are also based on the OpenRLHF framework under the Apache-2.0 license.
+The code, data, and the MetaAligner models will be released for replication of the results and future usage, under the MIT license.
+A.2 Broader Impacts
+In this work, MetaAligner provides an effective and model-agnostic method for generalizable and expandable alignment of LLM outputs with multiple human expectations.
+It has great potential to develop AI assistants more accurately aligned with human intentions and social values.
+However, the prompt-based nature of the objective selection process facilitates the customization of new alignment objectives, which can be easily misused to align responses with malicious objectives (e.g.
+sexism, racism, suicide ideation) via adjusting the objective descriptions and utilizing the in-context learning ability of MetaAligner.
+These actions can lead to harmful outputs from MetaAligner.
+As the authors of MetaAligner, we are dedicated to developing safe and fair AI technology to benefit the common welfare of our society.
+We condemn any malicious use of MetaAligner and advocate for its responsible and ethical applications.
+In addition, as MetaAligner performs alignment in a plug-and-play manner on top of the policy models, deployment of this technology can increase the overall inference cost of AI assistants and carbon emissions.
+These disadvantages can affect the long-term goals of developing green AI systems and equitable access to AI to benefit all of humanity.
+A.3 Safeguards
+This released codes, data, and MetaAligner models are provided for research only.
+None of the material constitutes actual diagnosis or advice, and help-seekers should get assistance from professional psychiatrists or clinical practitioners.
+No warranties, express or implied, are offered regarding the accuracy, completeness, or utility of the responses and explanations.
+The authors and contributors are not responsible for any errors, omissions, or any consequences arising from the use of the information herein.
+Users should exercise their own judgment and consult professionals before making any clinical-related decisions.
+The use of the software and information contained in this paper is entirely at the user’s own risk.
+The collected queries to build our IMHI preference dataset are from the publicly available IMHI dataset [34], and we strictly follow the privacy protocols and ethical principles to protect user privacy and guarantee that anonymity is properly applied in all the mental health-related texts.
+In addition, to minimize misuse, all examples provided in our paper are paraphrased and obfuscated utilizing the moderate disguising scheme.
+In addition, recent studies have indicated LLMs may introduce some potential bias, such as gender gaps.
+Meanwhile, some incorrect prediction results, inappropriate explanations, and over-generalization also illustrate the potential risks of current LLMs.
+Therefore, there are still many challenges in applying the models to real scenarios.
+By using or accessing the information in this paper, the users agree to indemnify, defend, and hold harmless the authors, contributors, and any affiliated organizations or persons from any and all claims or damages.
+Appendix B Text Descriptions and Prompting Templates
+The text descriptions for all tested objectives in this paper are included in Table 4.
+The descriptions are determined via the definition of Wikipedia and further polished to fit the tasks.
+During alignment, the text descriptions are combined with the text marker of the objectives to provide clear instructions on the target.
+The aligned objectives are obtained from the annotations of each alignment dataset, and the unaligned objectives are selected from the definitions of previous works [10].
+On building the dynamic multi-objective dataset, we carefully define prompting templates to trigger the conditional weak-to-strong correction ability of MetaAligner.
+A full list of the used templates is shown below.
+Specifically, for the preference subset of the HH-RLHF dataset [4], since the model is required to improve the response considering a multi-turn dialogue history, we have:
+[$\mathcal{T}(q,y,\mathcal{O},"better")$] You are an assistant to human.
+You will be provided with a context and an answer.
+Consider the context, then edit the answer to improve it considering these aspects: $\{\mathcal{O}\}$ | Context: {q} | Answer: {y} | Edit:
+For the equal-preference subset, we have:
+[$\mathcal{T}(q,y,\mathcal{O},"equal")$] You are an assistant to human.
+You will be provided with a context and an answer.
+Consider the context, then edit the answer to make it equal considering these aspects: $\mathcal{O}$} | Context: {q} | Answer: {y} | Edit:
+In the UltraFeedback dataset [8], the model is required to improve the response considering a single query. For the preference subset, we have:
+[$\mathcal{T}(q,y,\mathcal{O},"better")$] You are an assistant to human.
+You will be provided with a query and an answer.
+Consider the query, then edit the answer to improve it considering these aspects: $\mathcal{O}$} | Query: {q} | Answer: {y} | Edit:
+For the equal-preference subset, we have:
+[$\mathcal{T}(q,y,\mathcal{O},"equal")$] You are an assistant to human.
+You will be provided with a query and an answer.
+Consider the query, then edit the answer to make it equal considering these aspects: $\mathcal{O}$} | Query: {q} | Answer: {y} | Edit:
+Objectives $r$ Text Description $r\right>$
+Aligned Objectives
+Harmless The response should avoid content that is offensive, discriminatory, or harmful.
+Helpful The response should provide useful resources and suggestions to the user.
+Humor The response should be cheerful and amusing.
+Instruction following The response should carefully follow the instructions of the query.
+Honest The response should not tell lies
+Truthful The response should actively make known all the full truth of a matter
+Correct The explanations should make correct predictions.
+Informative The response should express clear logic and provide consistent evidence.
+Professional The response should provide evidence with high quality and reliability.
+Unaligned Objectives
+Specific The response should refer to facts and details and avoid vague arguments.
+Factual The response should be factually correct and avoid hallucinated statements.
+Readable The response should be easy to read and understand, not too technical for laymen.
+Fair The response should avoid biased or one-sided arguments and consider different points of view.
+Repeat The response should avoid repetitive statements of one point.
+Length The response should be concise and avoid redundant content.
+Table 4: Text descriptions for all tested objectives.
+Appendix C Experimental Details
+C.1 Model Training
+Details about the training process of MetaAligner and the building process of the 3 datasets are presented in Table 5.
+During the performance evaluation process, GPT-4 is leveraged to compare the responses on the corresponding objective.
+Specifically, we have the aligned test dataset: $\{q_{i},g_{i},R^{origin}_{i},R^{aligned}_{i}\}_{i=1}^{n}$, where $q$ denotes the query, $g$ denotes the ground-truth response from the dataset, $R^{origin}$ denotes the original response from the policy model, and $R^{aligned}$ denotes the aligned response from MetaAligner.
+We use the following prompting template and probe GPT-4 to perform judgment:
+[$\mathcal{E}(q,r_{1},r_{2},\left<r\right>)$] You are a skilled evaluator of helpful AI assistants.
+You will be presented with one query and two different responses to this query.
+QUERY: {q} | RESPONSE 1: {$r_{1}$} | RESPONSE 2: {$r_{2}$}.
+Consider the following aspect: $r\right>$}, then return the number of the better response.
+If tied, return 0.
+You must only return 1, 2, or 0.
+where $r_{1}$, $r_{2}$ are the compared response pairs, and $r\right>$ denotes the text description of the target objective.
+With the above information and the target objective description $r_{t}\right>$, we obtain the win rates using Algorithm 2.
+C.2 Model Cards
+TinyLLaMA-1.1B [36].
+A compact 1.1B language model pre-trained on around 1 trillion tokens for approximately 3 epochs.
+Building on the architecture and tokenizer of LLaMA2, TinyLlama leverages various advances contributed by the open-source community (e.g., Flash-Attention), achieving better computational efficiency.
+Despite its relatively small size, TinyLlama demonstrates remarkable performance in a series of downstream tasks.
+It significantly outperforms existing open-source language models with comparable sizes.
+We use TinyLlama-1.1B-Chat-v1.0 as the base model for MetaAligner-1.1B.
+LLaMA2-(Chat)-(7B, 13B, 70B) [31].
+A collection of pre-trained and fine-tuned large language models (LLMs) trained and released by Meta, ranging from 7 billion to 70 billion parameters.
+The fine-tuned LLMs, called LLaMA2-Chat, are optimized for dialogue use cases.
+The models outperform other open-source models on most benchmarks.
+Further human evaluations prove that LLaMA2-Chat also excels in helpfulness and safety.
+LLaMA2 models are among the most advanced open-source foundation models.
+We use LLaMA2-(7B, 13B) as base models for MetaAligner-(7B, 13B), and use LLaMA2-Chat-(7B, 13B, 70B) as policy models to evaluate the alignment performances.
+Vicuna-(7B, 13B, 33B) [7].
+Vicuna is a family of open-source chatbots trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT.
+Preliminary evaluation using GPT-4 as a judge shows Vicuna-13B achieves more than 90% quality of OpenAI ChatGPT and Google Bard while outperforming other models like LLaMA and Stanford Alpaca in more than 90% of cases.
+We use Vicuna-(7B, 13B)-V1.5 and Vicuna-33B-V1.3 as policy models to evaluate the alignment performances.
+Gemma-instruct-(2B, 7B) [30].
+A family of open-source models based on Google’s Gemini models.
+Gemma models are pretrained on 6T tokens of text, using architectures, data, and training recipes inspired by the Gemini model family.
+Like Gemini, these models achieve strong generalist capabilities in text domains, alongside state-of-the-art understanding and reasoning skills at scale.
+Gemma-instruct models are further fine-tuned for dialogue, instruction-following, helpfulness, and safety.
+Gemma-instruct is developed in two sizes: a 7B version for efficient deployment and development and a 2B version for CPU and on-device applications.
+We select both models as policy models to evaluate the alignment performances.
+MentaLLaMA-(7B, 13B, 33B) [34].
+MentaLLaMA is the first open-source instruction-following LLM series for interpretable mental health analysis.
+Based on LLaMA2-(7B, 13B) and Vicuna-33B foundation models, MentaLLaMA is trained on the Interpretable Mental Health Instruction (IMHI) dataset with 105K instruction samples, the first multi-task and multi-source instruction-tuning dataset for interpretable mental health analysis on social media.
+MentaLLaMA can perform mental health analysis on social media data and generate high-quality explanations for its predictions.
+On evaluating sub-task performance on IMHI Benchmark (Appendix D), we introduce MentaLLaMA-(7B, 13B, 33B) models as domain-specific policy models to evaluate the alignment performances.
+GPT-3.5 [22].
+GPT-3.5 is an advanced, close-sourced chat-based language model developed by OpenAI.
+It is a sibling model to InstructGPT, which is trained to follow instructions in a prompt and provide a detailed response.
+The model is firstly fine-tuned with SFT with conversations in which the model played both sides—the user and an AI assistant.
+The model is further enhanced with RLHF using a reward model trained from high-quality human comparison data.
+In our experiments, we use the gpt-3.5-turbo-0125 API provided by OpenAI as a strong policy model to evaluate the alignment performances.
+Claude-3 [3].
+Claude-3 is among the state-of-the-art foundation models for industry benchmarks across reasoning, math, coding, multi-lingual understanding, and vision quality, developed by Anthropic.
+The model family includes 3 models: (1) Opus, the most capable model; (2) Sonnet, which provides a combination of skills and speed; (3) Haiku, the fastest and least expensive model.
+All models are multi-modal and demonstrate strong performance across benchmark evaluations.
+Due to the budget limits, we select claude-3-sonnet-20240229 API provided by Anthropic as a strong policy model to evaluate the alignment performances.
+GPT-4 [2].
+Developed by OpenAI, GPT-4 is a large-scale, multimodal foundation model that can accept image and text inputs and produce text outputs.
+GPT-4 marks the highest level of achievement in AI industry and exhibits human-level performance on various professional and academic benchmarks, including passing a simulated bar exam with a score around the top 10% of test takers.
+We leverage the strong capability of GPT-4 and use it as an oracle to evaluate the large-scale test samples.
+Considering the high cost of evaluating large-scale test data and our limited budget, we use the cheaper GPT-4-turbo model with the gpt-4-turbo-preview API provided by OpenAI in practice.
+MetaAligner-(1.1B, 7B, 13B).
+Our proposed MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment.
+The models are based on TinyLLaMA and LLaMA2 foundation models.
+We train MetaAligner models on all 3 model scales for each of the 3 benchmark datasets.
+Specifically, HH-RLHF-MetaAligner is trained to align the responses of a general daily AI assistant with specified objectives considering multi-turn dialogue contexts.
+UltraFeedback-MetaAligner is trained to align responses of another general AI assistant considering a single-turn query, but the queries include professional questions such as programming language and history, and the aligned responses are usually more complicated.
+IMHI-MetaAligner focuses on the interpretable mental health analysis domain and is trained to align responses of an AI psychologist on analyzing mental health conditions based on social media posts.
+C.3 IMHI Annotation
+We select 1,200 queries from the IMHI benchmark covering 9 mental health analysis tasks.
+We obtain 4 responses to each query from 4 different policy models: GPT-4-turbo [2], GPT-3.5-turbo [22], MentaLLaMA-13B [34], and LLaMA2-Chat-13B [31], with human annotations on ranking different objectives of the responses.
+We utilize the above policy models to generate explanations for the same query simultaneously.
+The annotation protocol is developed through collaborative efforts with 2 domain experts (Ph.D.
+students majoring in quantitative psychology) and considerations of human evaluation criteria for previous mental health analysis tasks [34, 33].
+Specifically, 3 objectives are assessed: (1) Correctness: the explanations should make correct label predictions in the corresponding mental health analysis task; (2) Informativeness: the response should express clear logic and provide consistent evidence; (3) Professionalism: the response should provide evidence with high quality and reliability from the perspective of domain experts.
+Each aspect is divided into four standards rating from 0 to 3.
+Higher ratings reflect more satisfactory performance and 3 denotes approaching human performance.
+Each LLM-generated explanation is assigned a score by 2 domain experts for each corresponding objective, followed by the examination of 1 domain expert.
+All annotators are PhD students majoring in quantitative psychology.
+Annotators will be given generated responses from the 4 policy models and need to score and annotate the responses from the following objectives:
+Correctness.
+Correctness measures the trustworthiness of the classification results.
+Annotators should assess whether the classification result is based on facts, has misinformation, and wrong reasoning according to the given post.
+  • 0: Completely unreliable information with factual hallucination (e.g. non-existent symptoms).
+  • 1: Partly reliable information with wrong reasoning based on facts.
+  • 2: Mostly reliable information with non-critical misinformation or wrong reasoning.
+  • 3: Completely reliable information.
+Informativeness.
+Whether the text builds from sentence to sentence to a coherent body of information and logic about mental health and supports the classification results.
+Annotators should assess if the generated explanation gives consistent supporting evidence to its classifications and is well-structured.
+  • 0: Inconsistent with the classification results.
+  • 1: Consistent with the classification results, but with poor readability and several errors.
+  • 2: Consistent with the classification results. Mostly coherent and easy to read, with few minor errors.
+  • 3: Consistent with the classification results. Completely fluent, coherent, and error-free.
+Professionalism.
+Professionality measures the rationality of the generated explanations by evaluating the evidence that supports the classification results from the psychology perspective.
+Annotators should assess whether the explanation includes the following specified common diagnosis criteria of depression.
+To ensure the quality of the annotation scheme, we invite our domain experts to develop a list of common symptoms related to depression and sort these symptoms by criticality.
+The domain experts consult the Patient Health Questionnaire (PHQ-9) on determining the symptoms and sorting these symptoms on their knowledge.
+Specifically, the following symptoms are checked (sorted by criticality):
+  • Suicide ideation: Thoughts that you would be better off dead.
+  • Self-harm ideation: Thoughts of hurting yourself in some way.
+  • Feeling down, depressed, or hopeless.
+  • Self-guilt ideation: Feeling bad about yourself — or that you are a failure or have let yourself or your family down.
+  • Symptoms above are classified as with high criticality, and symptoms below are classified as with low criticality.
+  • Feeling tired or having little energy. Little interest or pleasure in doing things.
+  • Poor appetite or overeating.
+  • Trouble falling or staying asleep, or sleeping too much.
+  • Trouble concentrating on things, such as reading the newspaper or watching television.
+  • Moving or speaking so slowly that other people could have noticed. Or the opposite — being so fidgety or restless that you have been moving around a lot more than usual
+  • Uncontrollable sexual desire or sexual frigidity.
+  • Other symptoms.
+Based on the above symptoms, the annotators score the professionality of each explanation with the following criteria:
+  • 0: The explanation provides no supportive evidence or symptoms with high criticality are missing in the explanation.
+  • 1: The explanation provides a few supportive evidence, while some symptoms with higher criticality (than provided evidence) are missing.
+  • 2: The explanation provides several supportive evidence, while some symptoms with lower criticality (than provided evidence) are missing.
+  • 3: The explanation provides all related supportive evidence in the post.
+Algorithm 2 GPT-4 win-rate computation
+1:The aligned test dataset: $\{q_{i},g_{i},R^{origin}_{i},R^{aligned}_{i}\}_{i=1}^{n}$; Text description for target objective: $r_{t}\right>$; Prompting template: $\mathcal{E}(q,r_{1},r_{2},\left<r\right>)$
+2:Win rate of the aligned responses $\omega$.
+3:$W_{origin}\leftarrow\varnothing;W_{aligned}\leftarrow\varnothing$triangleright Initialize the judgement set W.
+4:$P_{origin}\leftarrow\varnothing;P_{aligned}\leftarrow\varnothing$triangleright Initialize the set P to record the position of the responses.
+5:$win_{origin}=0;win_{aligned}=0$triangleright Initialize the counter for wining samples.
+6:for $i\in\{1,...,n\}$ do
+7:     $r_{1}^{origin},r_{2}^{origin},p^{origin}_{i}=\textit{random\_shuffle}(g_{i},R^{origin}_{i})$triangleright Random shuffle the origin and ground-truth response.
+$p^{origin}_{i}$ denotes the position of $R^{origin}_{i}$.
+8:     $P_{origin}\leftarrow\mathcal{E}(q_{i},r_{1}^{origin},r_{2}^{origin},\left<r_{t}\right>)$triangleright Prompt for comparing origin and ground-truth response.
+9:     $J^{origin}_{i}\leftarrow\textit{Call-GPT-4}(P_{origin})$triangleright Call GPT-4 API to perform judgement.
+10:     $W_{origin}\leftarrow W_{origin}\cup\{J^{origin}_{i}\}$
+11:     $P_{origin}\leftarrow P_{origin}\cup\{p^{origin}_{i}\}$
+12:     if $J^{origin}_{i}=p^{origin}_{i}$ then
+13:         $win_{origin}=win_{origin}+1$
+14:     end if
+15:     $r_{1}^{aligned},r_{2}^{aligned},p^{aligned}_{i}=\textit{random\_shuffle}(g_{i},R^{aligned}_{i})$$\triangleright$ Similar actions for aligned response.
+16:     $P_{aligned}\leftarrow\mathcal{E}(q_{i},r_{1}^{aligned},r_{2}^{aligned},\left<r_{t}\right>)$
+17:     $J_{aligned}\leftarrow\textit{Call-GPT-4}(P_{aligned})$
+18:     $W_{aligned}\leftarrow W_{aligned}\cup\{J^{aligned}_{i}\}$
+19:     $P_{aligned}\leftarrow P_{aligned}\cup\{p^{aligned}_{i}\}$
+20:     if $J^{aligned}_{i}=p^{aligned}_{i}$ then
+21:         $win_{aligned}=win_{aligned}+1$
+22:     end if
+23:end for
+24:$\omega_{origin}=\frac{win_{origin}}{\textit{len}(W_{origin})}$triangleright Calculate win rates for original responses over ground-truth responses.
+25:$\omega_{aligned}=\frac{win_{aligned}}{\textit{len}(W_{aligned})}$triangleright Calculate win rates for aligned responses over ground-truth responses.
+26:$\omega=\omega_{aligned}-\omega_{origin}$
+Training Information
+Base Library Huggingface Transformers
+Fine-tuning Platform FastChat
+GPU Hardware 4$\times$ NVIDIA Tesla A100 80GB GPUs
+CPU Hardware 8$\times$ Intel(R) Xeon(R) Gold 6342 CPU cores per GPU
+Hardware Speedup Flash Attention 2 [9]
+Quantization for training BF16
+Fine-tuning Strategy Full fine-tuning
+Optimizer Adam
+Training Epochs 2
+Batch sizes HH-RLHF: 512 / UltraFeedback: 512 / IMHI: 128
+Max token for training MetaAligner-(1.1B, 7B, 13B): 2048/4096/4096
+Learning rate 1e-5
+Warm-up ratio 0.05
+Base Model-1.1B TinyLLaMA-1.1B
+Base Model-7B/13B LLaMA2-Chat-(7B, 13B)
+Dataset Information
+Dataset Name HH-RLHF
+Licence MIT
+Train/Val/Test ($\mathcal{D}_{p}$) 262,719/15,000/15,000
+Train/Val ($\mathcal{D}_{e}$) 16,502/1,797
+Harmless preference source Ray2333/gpt2-large-harmless-reward_model
+License MIT
+Helpful preference source Ray2333/gpt2-large-helpful-reward_model
+License MIT
+Humor preference source mohameddhiab/humor-no-humor
+License Apache-2.0
+Test evaluator GPT-4
+Dataset Name UltraFeedback
+Licence MIT
+Train/Val/Test ($\mathcal{D}_{p}$) 252,934/15,000/15,000
+Train/Val ($\mathcal{D}_{e}$) 82,023/5,000
+Instruction_following preference source GPT-4
+Honest preference source GPT-4
+Truthful preference source GPT-4
+Helpful preference source GPT-4
+Test evaluator GPT-4
+Dataset Name IMHI
+Licence MIT
+Train/Val/Test ($\mathcal{D}_{p}$) 5,304/1,051/2,400
+Train/Val ($\mathcal{D}_{e}$) 3,374/689
+Instruction_following preference source Human annotation
+Correct preference source Human annotation
+Informative preference source Human annotation
+Professional preference source Human annotation
+Test evaluator GPT-4
+Policy Models
+LLaMA2-Chat-(7B, 13B, 70B) https://huggingface.co/meta-llama
+Gemma-instruct-(2B, 7B) https://huggingface.co/google
+Vicuna-(7B, 13B, 33B) https://huggingface.co/lmsys
+GPT-3.5 https://openai.com/blog/chatgpt
+Claude-3-Sonnet https://www.anthropic.com/news/claude-3-family
+Table 5: Details for MetaAligner training and datasets. ‘preference source’ denotes how the preference annotations were obtained.
+Appendix D Sub-task Performance on IMHI Benchmark
+Data Task Source Annotation Labels/Aspects
+DR depression detection Reddit weak supervision Yes, No
+Dreaddit stress detection Reddit human annotation Yes, No
+SWMH mental disorders detection Reddit weak supervision Suicide, Anxiety, Bipolar disorder, Depression, None
+T-SID mental disorders detection Twitter weak supervision None, Suicide, Depression, PTSD
+SAD stress cause detection SMS human annotation
+School, Finance, Family, Social Relation,
+Work, Health, Emotion, Decision, Others
+CAMS depression/suicide cause detection Reddit human annotation
+Bias, Jobs, Medication, Relationship,
+Alienation, None
+loneliness loneliness detection Reddit human annotation Yes, No
+MultiWD Wellness dimensions detection Reddit human annotation
+Spiritual, Physical, Intellectual, Social,
+Vocational, Emotional
+IRF interpersonal risk factors detection Reddit human annotation Thwarted Belongingness, Perceived Burdensomeness
+Table 6: Details about the 9 sub-tasks in the IMHI dataset. "Annotation" denotes the reliability of the annotations in the raw data.
 MetaAligner Policy Model CAMS DR Dreaddit IRF loneliness MultiWD SAD SWMH T-SID Overall
@@
 Table 7: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models.
-The GPT-4 judge considers 3 objectives: Correctness, Informativeness, and Professionality.
+The GPT-4 judge considers 3 objectives: "Correct", "Informative", and "Professional".
 The figures show the advantage of aligned outputs over the policy model outputs on win rate.
 Best values for each MetaAligner model are highlighted in bold.
-4 Experiments
-4.1 Experimental Settings
-Datasets
-We train and evaluate the performance of MetaAligner on the IMHI dataset (Yang et al., 2023b).
-For the training data, we invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives with the following text markers: (1) Correctness: the response should make correct predictions in the corresponding sub-tasks; (2) informativeness: the response should express clear logic and provide consistent supporting evidence; (3) Professionality: the generated responses should provide supportive evidence with high quality and reliability.
-For evaluation, we perform alignment on the IMHI benchmark, which consists of 9 sub-tasks with 20K query-response pairs.
-Models
-We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B (Zhang et al., 2024) and LLaMA2-Chat-(7B, 13B) (Touvron et al., 2023) foundation models.
-We utilize MetaAligner to perform multi-objective alignment on the following policy models: LLaMA2-Chat-(7B,13B,70B) (Touvron et al., 2023), Gemma-instruct-(2B,7B) (Team et al., 2024), MentaLLaMA-(7B, 13B, 33B) (Yang et al., 2023b), and Vicuna-(7B, 13B, 33B) (Chiang et al., 2023).
-Among the policy models, MentaLLaMA models are instruction-tuned on mental health-related data, and other models are tuned on general-domain instruction data.
-Evaluation Metric
-We quantify the alignment performance of model outputs by comparing their win rate against the ground-truth response provided by the IMHI benchmark, on the corresponding objectives.
-Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works (Guo et al., 2024; Sun et al., 2023; Li et al., 2023a), to perform the judgments, where each target response, query, ground-truth response, and objectives are provided via prompt engineering.
-GPT-4 is required to select the response with higher alignment with the specified objectives in the prompt, or indicate a tied performance.
-Figure 2: Objective-wise win-rate comparisons based on different policy models.
-MetaAligner-7B with different levels of alignment are tested on Correctness, Informativeness, and Professionality, compared with 3 different policy models.
-The win rates are obtained via the judgments of GPT-4 on the corresponding objectives.
-4.2 Overall Performance
-We stack MetaAligner on different policy models to perform alignment on all 3 objectives: Correctness, Informativeness, and Professionality.
-The overall performance on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 2.
+We stack MetaAligner on different policy models to perform alignment on all 3 objectives: "Correct", "Informative", and "Professional".
+We include MentaLLaMA-(7B, 13B, 33B) [34], the first open-source instruction-following LLM series for interpretable mental health analysis into the policy models.
+Details about the 9 sub-tasks are provided in Table 6.
+The overall performance of MetaAligner on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 7.
 According to the results, the MetaAligner models achieve substantial improvement in overall performance on all 11 policy models, with an average of 26.89% advantage on win rates for MetaAligner-1.1B, 28.01% for MetaAligner-7B, and 36.6% for MetaAligner-13B.
@@
 These results show that MetaAligner can make reasonable corrections on weak responses while maintaining their expertise from domain-specific policy models.
-Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase.
-These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner.
-Fortunately, MetaAligner’s capabilities also show scalability.
-Increasing the size of its base model leads to a higher win-rate advantage on most policy models.
-Specifically, MetaAligner-7B outperforms or achieves comparable performance to MetaAligner-1.1B on 9 policy models, and MetaAligner-13B surpasses MetaAligner-7B on 10 policy models by an average of 12.58%.
-These observations motivate further explorations in model scale-performance balance for MetaAligner.
-Figure 3: Generalizable alignment on 3 unseen objectives based on LLaMA2-Chat-70B outputs.
-In the x-axis, "+" denotes the addition of a new objective into the objective set $\mathcal{O}$.
-"★" denotes the unaligned win rates for LLaMA2-Chat-70B outputs.
-"-." lines identify win rate fluctuations before alignment on the corresponding objective, and solid lines identify fluctuations after alignment.
-4.3 Objective-Wise Evaluation
-We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives.
-We utilize MetaAligner to perform 5 levels of alignments: No alignment, aligning on each objective (Correctness, Informativeness, and Professionality), and full alignment.
-The objective-wise win rates are shown in Figure 2.
-We have the following observations:
-Objective-wise alignment significantly improves the target objective and benefits the performance on other objectives.
-For example, aligning on Informativeness with MetaAligner achieves the best performance on Informativeness, and increases win rates on Correctness and Professionality over the unaligned outputs.
-This tendency holds with all policy models and other objectives.
-These results further prove the complex interplay within multi-objective alignment, where positive correlations also exist, in addition to contradictions (Guo et al., 2024).
-Aligning on all objectives provides balanced performance.
-According to the results, full alignment displays competitive performance on all 3 objectives and generally outperforms unaligned outputs and aligned outputs from other objectives, even comparable to aligned outputs from the same objective in some cases.
-The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, enabling the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
-MetaAligner performs comprehensive alignment on different sub-tasks.
-With most policy models, MetaAligner significantly outperforms the unaligned responses on all 9 sub-tasks, regardless of the alignment objectives.
-These results indicate the general effectiveness of our dynamic-objective method on tasks demanding various expertise (Yang et al., 2023b).
-However, MetaAligner still struggles to align certain sub-tasks, especially with strong policy models.
-For example, MetaAligner outputs fail to outperform MentaLLaMA-13B on the CAMS sub-task in all objectives.
-These results show enhancements in domain knowledge, with techniques such as retrieval-augment generation (Lewis et al., 2020), a promising future direction for improvement.
-Algorithm Space Complexity Correctness Informativeness Professionality Overall
-MORLHF 22.27$\times$ -6.31% +33.7% +16.31% +15.08%
-MODPO 17.27$\times$ +26.31% +30.1% +33.12% +32.9%
-SFT 11.82$\times$ +19.22% +24.89% +25.6% +26.62
-MetaAligner-1.1B 1$\times$ +26.11% +35.67% +13.11% +29.7%
-MetaAligner-7B 6.36$\times$ +28.89% +24.92% +29.07% +27.37%
-MetaAligner-13B 11.82$\times$ +31.33% +45.0% +43.66% +35.55%
-Table 3: Win rates of previous alignment methods and MetaAligner models based on the LLaMA2-Chat-13B policy model.
-Space complexity is calculated based on the memory consumption of training MetaAligner-1.1B.
-4.4 Generalizable Alignment
-In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 3 less relevant objectives that are unseen during its training phases: (1) Completeness (Yang et al., 2023a): the response should cover all relevant aspects of the original query; (2) Honesty (Guo et al., 2024): the response should speak in truth that is consistent with the query and commonsense.
-(3) Safety (Sun et al., 2023): The response should avoid content that is offensive, discriminatory, or harmful.
-2,700 random-selected responses from the IMHI benchmark are re-aligned with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one.
-Their performance is presented in Figure 3.
-According to the results, we have the following conclusions:
-MetaAligner performs effective zero-shot alignment for unseen objectives.
-With all MetaAligner models, incorporating an unseen objective significantly improves its corresponding win rate.
-For example, MetaAligner-7B improves win rate by 40.7% on Completeness, 28.89% on Honesty, and 17.56% on Safety compared to each of these objectives not introduced.
-These results prove the viability of generalizable alignment with the in-context learning ability of MetaAligner.
-In addition, alignment performance on supervised objectives (Correctness, Informativeness, Professionality) generally surpasses unseen objectives on all tested models, showing that MetaAligner also learns objective-specific features during training.
-Performance on aligned objectives is maintained with more unseen alignment objectives.
-As each objective is aligned, its win rate surges, then stabilizes as long as the objective is included in $\mathcal{O}$.
-These results prove MetaAligner to perform reliable alignment with the expansion of objectives.
-However, aligning on new objectives can affect other unaligned objectives.
-For example, aligning on Correctness with MetaAligner-1.1B benefits its win rate by 11.22%, but harms performance on 3 other unaligned objectives.
-These problems are due to the limited ability of smaller MetaAligner models and the controversial dependence of certain multi-objective alignments (Guo et al., 2024), where we observe an alleviation as the model size increases.
-MetaAligner’s generalizability also shows scalability.
-Performance on the 3 unseen objectives increases with the scale-up of MetaAligner model size, where MetaAligner-1.1B improves 12.7% on average, MetaAligner-7B improves 29.05%, and MetaAligner-13B improves 36.69%.
-This scalability is attributed to the growing in-context learning ability of larger base models, which motivates further explorations into the correlation between generalizable alignment and model scales.
-4.5 MetaAligner vs. MORLHF/MODPO/SFT
-We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on 2700 random-selected responses from the IMHI benchmark.
-We implement the linear scalarization method for MORLHF.
-We follow previous settings of realizing MODPO (Guo et al., 2024) by pre-pending the objective markers.
-Due to the closed-source nature of current SFT-based methods, we only replicate stage 1 where details are clear, and utilize the preference-to-reward mapping method proposed in Yang et al.
-(2024).
-We select LLaMA2-Chat-13B as the policy model, and the results are presented in Table 3.
-According to the results, MetaAligner-13B achieves over 30% of win rates on all objectives, and generally outperforms all previous methods by 10.68% on average.
-This advantage is achieved with equal or less space complexity than previous algorithms, showing the high efficiency of the proposed method.
-As the base model size further reduces, MetaAligner-1.1B and MetaAligner-7B show lower, but still competitive performance to the 13B model.
-Impressively, MetaAligner-1.1B outperforms both MORLHF and SFT-based methods by 8.85% on average, and approaches MODPO with merely a 3.2% gap, but with 11.82$\times$-22.27$\times$ less computational resources.
-This huge advantage indicates a high prospect for applying MetaAligner in low-resource scenarios.
-For previous methods, MODPO and SFT can provide stable alignment for multiple objectives, but with high space complexity.
-MORLHF bears the highest resource consumption, but achieves the lowest 15.08% overall win rate.
-Its training process is also unstable, including a 6.31% decrease in the Correctness objective.
-5 Related Work
-5.1 Large Language Models
-Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language Achiam et al.
-(2023); Team et al.
-(2023); Touvron et al.
-(2023); Luo et al.
-(2023).
-Notable examples include commercially available LLMs like ChatGPT, GPT-4 Achiam et al.
-(2023), and Gemini Team et al.
-(2023), as well as leading-edge open-source models like LLaMA2 Touvron et al.
-(2023) and Mistral Jiang et al.
-(2023).
-Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information Zhang et al.
-(2023); Xie et al.
-(2023), flattery, and deception, raising concerns about their potential negative impacts on society and associated risks Bommasani et al.
-(2021).
-To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning Ji et al.
-(2023).
-5.2 Preference Alignment
-Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty.
-These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) Ouyang et al.
-(2022).
-Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences.
-This limitation has motivated the exploration of multi-objective RLHF (MORLHF) Sener and Koltun (2018); Li et al.
-(2020); Rame et al.
-(2024), by linear scalarizations of multiple rewards Sener and Koltun (2018); Li et al.
-(2020) and interpolations of LLM weights trained from diverse reward models Rame et al.
-(2024).
-However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) Zhou et al.
-(2023) without the RL paradigm.
-Zhou et al.
-(2023) proposed multi-objective direct preference optimization (MODPO), which extended DPO algorithm (Rafailov et al., 2024) to combine multiple objectives with specific weightings.
-Another paradigm is the SFT-based methods Yang et al.
-(2024); Guo et al.
-(2024), which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO.
-These methods also facilitated objective-wise controllable generation during inference.
-6 Discussion
-Conclusion.
-This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment.
-Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives.
-Thorough investigations on LLMs ranging from 2B to 70B proved MetaAligner’s overall effectiveness in multi-objective alignment for various policy models and objective-wise alignment.
-and Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align more objectives.
-Limitation.
-Though achieving outstanding performance, MetaAligner still bears limitations.
-Firstly, similar to the vanilla Aligner module Ji et al.
-(2024), stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference, which affects actual model deployment, especially for scenarios such as local deployment on mobile devices.
-Secondly, MetaAligner still provides limited improvement for large-scale domain-specific policy models, such as MentaLLaMA-33B, in domain knowledge-intense tasks such as mental health analysis.
-These results reflect the importance of expertise knowledge for multi-objective alignment, where MetaAligner still requires further enhancements.
-Thirdly, due to limited resources, we only tested the generalizability of MetaAligner on 3 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives.
-Future Work.
-Firstly, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation (Lewis et al., 2020).
-Secondly, we will dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance.
-We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 6 objectives).
-This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
+Appendix E Objective-wise Alignment
+E.1 Experimental Settings
+We randomly sample 1,200 queries from the UltraFeedback test set and probe the target policy models to provide responses to all queries, which are regarded as unaligned responses.
+Then MetaAligner is used to align these responses under different objectives, including aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and another full alignment on all 4 objectives.
+During evaluation, GPT-4 is leveraged as a reward model to score these responses considering different objectives.
+Specifically, the following prompt is developed to obtain the scores:
+[$\mathcal{S}(q,a,\left<r\right>)$] You are a skilled evaluator of helpful AI assistants.
+You will be presented with one query and a response to this query.
+QUERY: {q} | RESPONSE: {a} Assign a score ranging from 0 to 10 to this response considering the following aspect: $r\right>$}.
+The rating improves as the score rises, where 0 denotes inferior performance and 10 denotes approaching-human performance.
+You must only return the assigned score.
+where $q$ and $a$ denotes the target query and response, and $r\in${Instruction following, Helpful, Honest, Truthful} is the target objective for evaluation.
+E.2 Experimental Results
+More experimental results are presented in Figure 4.
+The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5 outputs from the UltraFeedback test set.
+According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5.
+Unlike full alignment which shows weaker performance as the capability of the policy model increases, objective-wise alignment provides stable improvement for different policy models.
+For example, aligning "Instruction following" leads to over 20% of approaching-human responses for all policy models, and aligning "Honesty" leads to over 30% of approaching-human responses for all policy models.
+The reason is that single-objective alignment doesn’t involve the complex interactions between multiple objectives, which facilitates the full realization of MetaAligner’s potential to improve the response on the corresponding objective.
+Figure 4: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
+Appendix F Implementation Details of Baseline Models
+F.1 SFT-based Methods
+Existing SFT-based methods [11, 35] for multi-objective preference alignment explicitly incorporate reward values into the query via prompt engineering, which includes a text marker for each alignment objective and their corresponding value numbers in the current response.
+The model is trained to predict the response given the enhanced query, which enables it to learn a mapping between the reward value and its reflection in the response.
+During inference, we achieve alignment by assigning a higher reward value to the target objectives.
+Specifically, we define the following prompting template for HH-RLHF dataset:
+[$\mathcal{F}(q,r_{1},r_{2},r_{3}$] <Harmlessness>: {$r_{1}$}; <Helpfulness>: {$r_{2}$}; <Humour>: {$r_{3}$} | {q}
+where $q$ denotes the query, $r_{1}$, $r_{2}$, $r_{3}$ denote the corresponding reward values for the current response, ranging from 1 to 5.
+We define the following prompting template for UltraFeedback dataset:
+[$\mathcal{F}(q,r_{1},r_{2},r_{3},r_{4}$] <instruction_following>: {$r_{1}$}; <honesty>: {$r_{2}$}; <honesty>: {$r_{3}$}; <helpfulness>: {$r_{4}$} | {q}
+where $q$ denotes the query, $r_{1}$, $r_{2}$, $r_{3}$, $r_{4}$ denote the corresponding reward values for the current response, obtained from existing reward models.
+During inference, we aim to simultaneously align all objectives with 1 model to enable fair comparisons to MetaAligner.
+For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during inference: $r_{1}=r_{2}=r_{3}=r_{4}=5$.
+For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models.
+Therefore, we set each objective value to its maximum in the training dataset to enable higher alignment.
+The values are set to: $r_{1}=4.19$, $r_{2}=3.03$, $r_{1}=1.58$.
+In calculating the GPU hours, we sum the training hours for fine-tuning the LLaMA-Chat-7B policy model for HH-RLHF and UltraFeedback datasets.
+F.2 MODPO
+We implement MODPO by implementing the Controllable Direct Preference Optimization (CDPO) [11] algorithm based on its paper introduction.
+We bypass the controllable preference SFT stage by utilizing the trained model from SFT-based methods as the foundation model.
+In the CDPO stage, each query $q$ is accompanied by two responses $c_{1}$ and $c_{2}$, where each response is constructed into contrastive pairs.
+In HH-RLHF, we have the following prompting template:
+[$\mathcal{F}(q,c_{1},r_{1}^{1},r_{1}^{2},r_{1}^{3}$] <Harmlessness>: {$r_{1}^{1}$}; <Helpfulness>: {$r_{1}^{2}$}; <Humour>: {$r_{1}^{3}$} | {q} | {$c_{1}$}
+[$\mathcal{F}(q,c_{2},r_{2}^{1},r_{2}^{2},r_{2}^{3}$] <Harmlessness>: {$r_{2}^{1}$}; <Helpfulness>: {$r_{2}^{2}$}; <Humour>: {$r_{2}^{3}$} | {q} | {$c_{2}$}
+where $r_{i}^{j}$ denotes the reward value for the $i$-th response on the $j$-th objective.
+For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models.
+Therefore, we set each preference value to its maximum in the training dataset to enable higher alignment.
+The values are set to: $r^{1}=4.19$, $r^{2}=3.03$, $r^{3}=1.58$.
+Similarly, on UltraFeedback we have:
+[$\mathcal{F}(q,c_{1},r_{1}^{1},r_{1}^{2},r_{1}^{3},r_{1}^{4}$] <instruction_following>: {$r_{1}^{1}$}; <honesty>: {$r_{1}^{2}$}; <honesty>: {$r_{1}^{3}$}; <helpfulness>: {$r_{1}^{4}$} | {q} | {$c_{1}$}
+[$\mathcal{F}(q,c_{2},r_{2}^{1},r_{2}^{2},r_{2}^{3},r_{2}^{4}$] <instruction_following>: {$r_{2}^{1}$}; <honesty>: {$r_{2}^{2}$}; <honesty>: {$r_{2}^{3}$}; <helpfulness>: {$r_{2}^{4}$} | {q} | {$c_{2}$}
+For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during the DPO training process: $r^{1}=r^{2}=r^{3}=r^{4}=5$.
+In determining the preference scores, we use the CDPO learning goal to transform the task into a conditional multi-objective optimization problem.
+Specifically, the reward value $R_{i}$ for response $c_{i}$ is calculated as follows:
+$R_{i}=\sum_{i=1}^{m}\omega_{i}g_{i}$ (9)
+where $\omega_{i}$ represents the weight of the $i$-th objective and $g_{i}$ is calculated as follows:
+$\small g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (10)
+where $\lambda_{i}$ represents the weight of the controlled objective and $p_{j}$ is a pre-defined preference value for the $j$-th objective.
+In our implementation, we set $\omega_{i}=\lambda_{i}=1$.
+We aim to simultaneously align all objectives with 1 model to enable fair comparisons to MetaAligner.
+For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during inference: $p_{1}=p_{2}=p_{3}=p_{4}=5$.
+For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models.
+Therefore, we set each preference value to its maximum in the training dataset to enable higher alignment.
+The values are set to: $p_{1}=4.19$, $p_{2}=3.03$, $p_{3}=1.58$.
+After calculation, the response with a higher reward value $R_{i}$ is used as the chosen response, and the other response is used as the rejected response for MODPO training.
+Specifically, the model is trained via the following loss function:
+$\mathcal{L}_{\mathrm{CDPO}}={\tiny-\mathbb{E}_{\left(x,c,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid c,x)}{\pi_{\text{ref }}(y_{w}\mid c,x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid c,x)}{\pi_{\text{ref }}(y_{l}\mid c,x)}\right)\right]}$ (11)
+where $x$ denotes the query, $y_{w},y_{l}$ denote the chosen and rejected prompts, $c$ denotes the corresponding condition, $\pi_{\theta}$ and $\pi_{ref}$ denote the target policy model and the reference model.
+For implementation, we build our MODPO code based on the OpenRLHF library.
+In calculating the GPU hours, we include the training hours for tuning the SFT-based policy model using the CDPO algorithm for HH-RLHF and UltraFeedback datasets.
+We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for fair comparisons.
+F.3 MORLHF
+We use the linear scalarization [27, 20] realization of MORLHF with the KL-divergence regularization. Specifically, the model is trained via the following objective function:
+$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{ref}(y|q)}\right]$ (12)
+where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector, $q$, $y$ denote the query and the response, R denotes the reward functions for the target objectives.
+In implementing the MORLHF algorithm, we first train a reward model for each objective based on random samples from 50% of the training data.
+Following the reward models we used in building the HH-RLHF dataset, we select the GPT2-large as foundation models for the reward model, and optimize the reward models using the following pair-wise loss functions:
+$\mathcal{L}_{rm}=-log(\sigma(R_{c}-R_{r}-margin))$ (13)
+where $\sigma$ denotes the Sigmoid function, $R_{c}$ and $R_{r}$ denote the reward output of the chosen response and the rejected response, and $margin$ denotes the margin loss for the corresponding response pairs when multiple responses are ranked, such as in UltraFeedback.
+Secondly, we fine-tune the LLaMA2-Chat-7B policy model with the highest-ranked responses from the HH-RLHF and UltraFeedback datasets to obtain sub-optimal starting points for RLHF.
+The SFT training process is formalized as follows:
+$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q)\right]$ (14)
+where $q$ and $y$ denote the query and its corresponding highest-ranked response.
+Thirdly, following most works in RLHF, we leverage the PPO algorithm [26] to enable parameterized optimization of the policy model.
+In linear scalarization, we set $\omega_{1}=\omega_{2}=...=\omega_{N}=\frac{1}{N}$.
+For implementation, we build our MORLHF code based on the OpenRLHF library.
+In calculating the GPU hours, we include the training hours for all reward models in HH-RLHF and UltraFeedback datasets, with a sum of seven reward model training processes.
+We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for reaching the sub-optimal starting points.
+Finally, the PPO training hours for HH-RLHF and UltraFeedback are included in the GPU hours.

### v3

@@
 Abstract
-Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment.
-However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives.
-In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses.
-MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning.
-Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and outperforms previous alignment methods with down to 15.71$\times$ less GPU training hours.
+Recent advancements in large language models (LLMs) focus on aligning to heterogeneous human expectations and values via multi-objective preference alignment.
+However, existing methods are dependent on the policy model parameters, which require high-cost repetition of their alignment algorithms for each new policy model, and they cannot expand to unseen objectives due to their static alignment objectives.
+In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment.
+MetaAligner models multi-objective alignment into three stages: (1) dynamic objectives reformulation algorithm reorganizes traditional alignment datasets to supervise the model on performing flexible alignment across different objectives; (2) conditional weak-to-strong correction paradigm aligns the weak outputs of fixed policy models to approach strong outputs with higher preferences in the corresponding alignment objectives, enabling plug-and-play inferences on any policy models, which significantly reduces training costs and facilitates alignment on close-source policy models; (3) generalizable inference method flexibly adjusts target objectives by updating their text descriptions in the prompts, facilitating generalizable alignment to unseen objectives.
+Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and saves up to 93.63% of GPU training hours compared to previous alignment methods.
 The model also effectively aligns unseen objectives, marking the first step towards generalizable multi-objective preference alignment.
-This project is open-sourced at: https://github.com/SteveKGYang/MetaAligner
+This project is open-sourced here.
 1 Introduction
-The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values.
-At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [39, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model [23].
+The recent advancements in large language models (LLMs) have focused on generating high-quality responses that align with human expectations and values.
+At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [41, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning (RL) on the target model [23].
 However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences [4, 25].
@@
 Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) [27, 20, 25].
-Due to its substantial computational cost [20, 25] and the unstable nature of the PPO [26, 17, 24], other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [38, 11] or supervised fine-tuning (SFT)-based methods [35, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
-The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model.
-This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g.
+Due to its substantial computational cost [20, 25] and the unstable nature of the proximal policy optimization (PPO) [26, 16, 24] algorithm, other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [40, 11] or supervised fine-tuning (SFT)-based methods [36, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
+The above methods for multi-objective alignment bear one commonality: the dependence on the policy model’s parameters.
+This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced policy model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g.
 "Helpful", "Harmless", "Honest" [40, 11]) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives.
 This ignorance leads to poor generalizability of existing multi-objective alignment methods.
-Algorithm Paradigm Reward-Free Align. Multi-Objective Align. Policy-Agnostic Align. Generalizability
+Table 1: Comparisons between previous alignment methods and MetaAligner on different features.
+"Policy-Agnostic Alignment" means the alignment algorithm is independent of the target policy model parameters, and "Generalizability" denotes zero-shot alignment capability on unseen objectives.
+Algorithm Paradigm Multi-Objective Alignment Policy-Agnostic Alignment Generalizability
 RLHF [23] PPO
@@
 MetaAligner SFT
-Table 1: A checklist for key characteristics of previous preference alignment methods and MetaAligner. "Align." denotes "Alignment".
 In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment.
-MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access [13].
-Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives.
-The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction.
-That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering.
-During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections.
-As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment.
-The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning [16].
-The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives.
-A checklist of MetaAligner’s advantages over previous methods is presented in Table 1.
-Through this work, we make three main contributions: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment.
+MetaAligner models multi-objective alignment into three stages: (1) the dynamic objectives reformulation algorithm reorganizes traditional alignment datasets into dynamic-objective alignment datasets, training MetaAligner to perform flexible alignment across different objectives.
+It achieves this by incorporating and combining text descriptions of various alignment objectives in a prompt-based manner; (2) the conditional weak-to-strong correction paradigm aligns the weak outputs of policy models to approach strong outputs with higher preferences in the corresponding alignment objectives.
+During training, MetaAligner is stacked onto policy models to perform objective-aware corrections, where parameters of the policy model are fixed and MetaAligner is optimized with an SFT-based three-step training process: warming up, equal-preference alignment, and contrastive-preference alignment.
+This paradigm enables MetaAligner to perform plug-and-play inferences on any policy models even without access to their parameters, which significantly reduces training costs and facilitates alignment on close-source LLMs; (3) the generalizable inference method flexibly adjusts target objectives by updating their text descriptions in the prompts.
+This method can also adapt MetaAligner to unseen objectives and achieve new alignment strategies via in-context learning [15], a new feature with rare previous exploration in alignment of language models.
+The number of aligned objectives also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives.
+Table 1 compares key features between MetaAligner and previous methods.
+As shown, conditional weak-to-strong correction of MetaAligner extends Aligner [13] to multi-objective alignment scenarios, which are not directly solvable by Aligner itself.
+MetaAligner is also the first multi-objective alignment method to achieve policy-agnostic alignment and generalization to unseen objectives, two key advantages over previous methods such as MORLHF, MODPO, and SFT-based methods.
+In summary, our main contributions are: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment.
 It performs multi-objective alignment efficiently, without tuning the policy models or accessing their parameters.
-Experimental results show that MetaAligner outperforms previous alignment methods with down to 15.71× less GPU training hours and higher stability; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives.
+Experimental results show that MetaAligner outperforms previous alignment methods and saves up to 93.63% of GPU training hours; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives.
 To our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment.
-Experimental results show that MetaAligner can simultaneously perform effective alignment for 6 unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on 3 preference alignment datasets.
-Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5 [22] and Claude-3 [3].
-Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF [20], (B) SFT-based methods [35, 11], and (C) Meta-Objective Aligner.
-SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
+Experimental results show that MetaAligner can simultaneously perform effective alignment for six unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on three preference alignment datasets.
+Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5-Turbo [22] and Claude-3 [3].
+Figure 1: Illustrations of Meta-Objective Aligner, which follows a three-stage paradigm.
 2 Multi-Objective Alignment of Language Models
 In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplays such as contradiction (e.g.
-"Helpful" and "Harmless") and dependence (e.g.
-"Correct" and "Informative").
-Multi-objective preference alignment tackles this challenge by optimizing multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$.
-$r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences.
+"Helpful" and "Harmless") [11, 36] and dependence (e.g.
+"Correct" and "Informative") [35].
+Multi-objective preference alignment tackles this challenge by aiming to optimize multiple objectives simultaneously.
+For each query-response pair, the reward vector is formalized as: $\mathbf{R}(q,y)=[r_{1}(q,y),...,r_{N}(q,y)]^{\mathrm{T}}$, where $q,y$ denote a query and a corresponding response, $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences.
 Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
-$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\frac{1}{1+exp(-(r_{i}(q,y_{1})-r_{i}(q,y_{2})))}$ (1)
-where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$.
-MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization (see Figure 1 (A)):
-$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
-where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector.
-Another paradigm [35, 11] directly built alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
-$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q,\mathbf{R}(q,y))\right]$ (3)
+$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\sigma(r_{i}(q,y_{1})-r_{i}(q,y_{2}))$ (1)
+where $\sigma$ denotes the logistic function and $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$.
+MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization:
+$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}(y|q)}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)\right]-\beta\mathbb{D}_{KL}\left[\pi_{\phi}(y|q)\|\pi_{ref}(y|q)\right]$ (2)
+where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$, $\pi_{ref}$ denotes the reference policy model, $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the pre-determined heuristic target preference vector.
+Another paradigm directly built alignment between multiple reward values and their corresponding response by minimizing an SFT loss for the policy model:
+$\mathop{argmin}\limits_{\pi_{\phi}}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[\mathop{log}\pi_{\phi}(y|q,\mathbf{R}(q,y))\right]$ (3)
 where objectives and their corresponding reward values are described with text markers and combined into queries with a static prompting template.
-Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable.
+Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable [36, 11].
 3 Meta-Objective Aligner
-This section introduces the proposed methodology for MetaAligner, which follows a 3-stage paradigm: (1) dynamic multi-objective dataset construction; (2) a 3-step model training process; (3) generalizable inference for alignment.
-The paradigm is illustrated in Figure 1 (C).
-3.1 Dynamic Multi-Objective Dataset
+Existing methods for multi-objective alignment generally face challenges in increasing training costs with new policy models and generalization to unseen objectives.
+To tackle these challenges, we introduce MetaAligner, which follows a three-stage paradigm: (1) dynamic objectives reformulation for building dynamic multi-objective datasets; (2) conditional weak-to-strong correction for model training; (3) generalizable inference for multi-objective alignment.
+The paradigm is illustrated in Figure 1.
+3.1 Dynamic Objectives Reformulation
+We propose a dynamic objectives reformulation algorithm to construct a dynamic multi-objective dataset, which triggers MetaAligner’s ability for flexible adjustment of alignment objectives.
+Specifically, any typical multi-objective preference alignment dataset $\mathcal{D}_{m}$ with $m$ samples and $N$ objectives can be re-organized as $\{q_{i},y_{i1},y_{i2},P_{i}\}_{i=1}^{m}$, where $P_{i}=[p_{i1},...,p_{iN}]^{\mathrm{T}}$ and $p_{ij}\in\{\succ,\prec,\equiv\}$ indicates the preference on $j$-th objective.
+Algorithm 1 Dynamic objectives reformulation.
+[…]
+2:Contrastive subset $\mathcal{D}_{c}$; Equal subset $\mathcal{D}_{e}$.
+3:$\mathcal{D}_{c}\leftarrow\varnothing$, $\mathcal{D}_{e}\leftarrow\varnothing$triangleright Initialize the 2 subsets.
+4:for $i\in\{1,...,m\}$ do $\triangleright$ Loop on instances.
+5:     $\mathcal{O}_{\succ}\leftarrow\varnothing$, $\mathcal{O}_{\prec}\leftarrow\varnothing$, $\mathcal{O}_{\equiv}\leftarrow\varnothing$
+6:     for $j\in\{1,...,N\}$ do
+7:         if $p_{ij}$ is $\succ$ then $\triangleright$ Collect the objectives where $y_{i1}$ outperforms $y_{i2}$.
+8:              $\mathcal{O}_{\succ}\leftarrow\mathcal{O}_{\succ}\cup\{\left<d_{j}\right>\}$
+9:         else if $p_{ij}$ is $\prec$ $\triangleright$ Collect the objectives where $y_{i2}$ outperforms $y_{i1}$.
+10:              $\mathcal{O}_{\prec}\leftarrow\mathcal{O}_{\prec}\cup\{\left<d_{j}\right>\}$
+11:         $\triangleright$ Collect the objectives where $y_{1}$ and $y_{2}$ performs equally.
+12:              $\mathcal{O}_{\equiv}\leftarrow\mathcal{O}_{\equiv}\cup\{\left<d_{j}\right>\}$
+13:         end if
+14:     end for
+15:     if $\mathcal{O}_{\succ}\neq\varnothing$ $\triangleright$ Build the training pairs where $y_{i1}$ is used as the target.
+16:         $t\leftarrow better$
+17:         $\mathcal{O}_{\succ}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\succ})$
+18:         $\mathcal{D}_{c}\leftarrow\mathcal{D}_{c}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\succ},t),y_{i1})\}$
+19:     end if
+20:     if $\mathcal{O}_{\prec}\neq\varnothing$ $\triangleright$ Build the training pairs where $y_{i2}$ is used as the target.
+21:         $t\leftarrow better$
+22:         $\mathcal{O}_{\prec}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\prec})$
+23:         $\mathcal{D}_{c}\leftarrow\mathcal{D}_{c}\cup\{(\mathcal{T}(q_{i},y_{i1},\mathcal{O}_{\prec},t),y_{i2})\}$
+24:     end if
+25:     if $\mathcal{O}_{\equiv}\neq\varnothing$ $\triangleright$ Build equally-preferred training pairs.
+26:         $t\leftarrow equal$
+27:         $\mathcal{O}_{\equiv}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\equiv})$
+28:         $\mathcal{D}_{e}\leftarrow\mathcal{D}_{e}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\equiv},t),y_{i1})\}$
+29:     end if
+30:end for
 […]
 [$\mathcal{T}(q,y,\mathcal{O},t)$] Edit the following Question-Answer pair to make it {$t$} considering the following objectives $\mathcal{O}$} | Question: {$q$} | Answer: {$y$} | Edit:
-where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the building subset.
-The detailed construction process is described in Algorithm 1.
-A full list of the utilized prompting templates is presented in Appendix B.
-Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping [35] process in previous SFT-based multi-objective alignment methods.
-3.2 MetaAligner Derivation
-Based on the dynamic multi-objective dataset, we achieve multi-objective alignment by introducing MetaAligner, a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the initial outputs as follows:
+where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the concatenation of text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the current building subset.
+Details of the dynamic objectives reformulation algorithm are described in Algorithm 1.
+For an instance within the processing loop (line 2): $\{q,y_{1},y_{2},P\}$, the algorithm performs a two-step reformulation: (1) collect the sets $\mathcal{O}_{\succ}$, $\mathcal{O}_{\prec}$, $\mathcal{O}_{\equiv}$ that includes objectives where $y_{1}$ outperforms $y_{2}$, $y_{2}$ outperforms $y_{1}$, and both perform equally (lines 3-12); (2) for each objective set $\mathcal{O}$, we randomly shuffle the objectives to further trigger the model’s flexible alignment ability, and build query-response pairs based on the corresponding prompting template $\mathcal{T}$ (line 13-27).
+All prompting templates and examples of the algorithm are presented in Appendix C.
+Training on dynamic multi-objective datasets provides three key advantages: (1) instance-level alternation of the target objectives during training enables MetaAligner to perform flexible alignment under different combinations of objectives; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors.
+(3) the reward-free alignment method (no explicit preference values required) avoids complicated preference-to-reward mapping [36] process in previous SFT-based multi-objective alignment methods.
+3.2 Conditional Weak-to-Strong Correction
+Based on the dynamic multi-objective training datasets, we train MetaAligner in a conditional weak-to-strong correction manner, which follows an SFT-based training objective and a three-step training paradigm.
+3.2.1 Training Objective Derivation
+MetaAligner is a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the policy model output $y_{0}$ considering objectives $\mathcal{O}$ as follows:
 $\pi^{*}(y|q)=\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\pi_{\phi}(y_{0}|q)$ (4)
-where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$.
-We achieve alignment by directly employing SFT on the target response.
+where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$, $t$ depends on the training dataset.
+Conditional weak-to-strong correction directly trains MetaAligner to align the weak policy model output $y_{0}$ to the strong target output $y$, which has higher preference values in corresponding objectives $\mathcal{O}$.
 We have the standard cross-entropy loss as the training objective:
-$\mathop{argmin}\limits_{\theta,\phi}$ $\mathcal{L}(\theta,\phi;\mathcal{D})=-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\pi^{*}(y|q)\right]$ (5)
-=-= - $\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-$
-$\mathbb{E}_{q\sim\mathcal{D}}\left[log\pi_{\phi}(y_{0}|q)\right]$
-We fuse the parameters of the policy model, thus excluding $\phi$ from the weight update process.
-As only $\theta$ derivatives are calculated during back-propagation, we eliminate the second term in Eqn 5 and simplify the training objective as:
-$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
-The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as samples from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs [2, 22, 3].
-3.3 Model Training
+$\mathop{argmin}\limits_{\theta,\phi}\mathcal{L}(\theta,\phi;\mathcal{D})$ $=-\mathbb{E}_{(q,y,\mathcal{O})\sim\mathcal{D}}\left[\mathop{log}\pi^{*}(y|q)\right]$ (5)
+$=-\mathbb{E}_{(q,y,\mathcal{O})\sim\mathcal{D}}\left[\mathop{log}\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-\mathbb{E}_{q\sim\mathcal{D}}\left[\mathop{log}\pi_{\phi}(y_{0}|q)\right]$
+We fix the parameters of the policy model, thus excluding $\phi$ from the weight update process.
+In practice, we use the dynamic multi-objective dataset for supervision, where the weak response in each query-response pair is directly leveraged as samples $y_{0}$ from unknown policy models.
+Therefore, we eliminate the second term in Eqn.
+5 and simplify the training objective as:
+$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y_{0},y,\mathcal{O})\sim\mathcal{D}}\left[\mathop{log}\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
+The above action poses two advantages: (1) the computation resources required for MetaAligner training is detached from policy model size, which enables policy-agnostic and cost-efficient alignment for large policy models; (2) MetaAligner works only via outputs from the policy models, which allows training and inference for alignment on close-source policy models [2, 22, 3].
+3.2.2 Three-Step Model Training
 In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process.
-We propose a 3-step paradigm based on the objective function in Eqn.
-6:
-Step 1: warm-up.
-The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies [12, 13] to familiarize the model with identity mapping and facilitate convergence.
-Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
-Step 2: equal-preference modeling.
+We propose a three-step paradigm based on the objective function in Eqn.
+6: (1) Warming up.
+This stage trains the model in identical response pairs with a warm-up subset, a prelude proven effective in residual correction strategies [12, 13].
+We randomly sample a subset of the equal subset $\mathcal{D}_{e}$ as the warm-up subset, but set an identical target response for each instance; (2) Equal-preference alignment.
 Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs.
 However, equal preferences are common in many scenarios [35, 8] and enclose useful information such as the principle components of preference modeling regarding each objective.
-Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
-Step 3: preference alignment
-This stage fine-tunes the MetaAligner on the preference subset $\mathcal{D}_{p}$, which instructs the model to perform conditional weak-to-strong correction on the dis-preferred response and achieve alignment to the preferred response on the specified objectives.
+Based on these intuitions, we introduce a novel equal-preference alignment step to fine-tune the warmed-up MetaAligner on the equal subset $\mathcal{D}_{e}$; (3) Contrastive-preference alignment.
+This stage fine-tunes the MetaAligner on the contrastive preference subset $\mathcal{D}_{c}$, which instructs the model to perform conditional weak-to-strong correction on the specified objectives.
 3.3 Generalizable Inference
 […]
-[…]
-[…]
-[…]
-This simple pattern can theoretically lead to unlimited simultaneous alignment objectives.
-We expect MetaAligner to make generalizable weak-to-strong corrections under these unseen conditions via its in-context learning ability.
-This advancement marks a new exploration into generalizable multi-objective preference alignment.
 4 Experiments
@@
 Datasets.
-We train and evaluate on 3 alignment datasets: (1) HH-RLHF [4]: a large-scale dataset with 160K prompts and corresponding response pairs.
-We follow previous works [35] and utilize 3 open-sourced reward models on "Harmless", "Helpful", and "Humor" to assess and rank the responses in their corresponding objectives; (2) UltraFeedback [8]: a multi-aspect alignment dataset with 64K prompts with preferences obtained from GPT-4.
-We select "Instruction following", "Honest", "Truthful", and "Helpful" objectives during training; (3) IMHI: we evaluate MetaAligner on domain knowledge-intense scenarios by creating another alignment dataset based on the IMHI dataset [34] targeting interpretable mental health analysis, which consists of 9 mental health analysis sub-tasks.
-We invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives: "Correct", "Informative", and "Professional".
+We transfer the following three alignment datasets into dynamic multi-objective datasets: (1) HH-RLHF [4]: a large-scale dataset with 160K prompts and corresponding response pairs.
+We follow Yang et al.
+[36] and use open-sourced reward models on three objectives: "Harmless", "Helpful", and "Humor" to score the responses; (2) UltraFeedback [8]: a multi-aspect alignment dataset with 64K prompts with preferences obtained from GPT-4, including "Instruction following", "Honest", "Truthful", and "Helpful" objectives; (3) IMHI: we create an alignment dataset on the IMHI dataset [35] targeting interpretable mental health analysis.
+We invite domain experts to label 7.2K response pairs considering 3 objectives: "Correct", "Informative", and "Professional".
+Figure 2 shows the objective distributions on two datasets.
+The objectives display balanced overall distributions across objective set sizes, training MetaAligner to adjust targets dynamically.
+Most objectives also cover considerable proportions in each column category, alleviating label imbalance problems.
+Figure 2: Heatmaps of the objective distributions.
+The columns categorize samples according to the sizes of their objective set.
+For the lines, "Overall" shows their distributions in the training data.
+Other lines show objective-wise distributions across different categories in the columns.
 Models.
@@
 We utilize MetaAligner to perform multi-objective alignment on the following open-source policy models: LLaMA2-Chat-(7B,13B,70B) [31], Gemma-instruct-(2B,7B) [30], and Vicuna-(7B, 13B, 33B) [7].
-We also align 2 advanced close-sourced foundation models: GPT-3.5 [22] and Claude-3-Sonnet [3], where model parameters are inaccessible.
+We also align two advanced close-source foundation models: GPT-3.5-Turbo [22] and Claude-3-Sonnet [3], where model parameters are inaccessible.
 Evaluation Metric.
 On each objective, we quantify the alignment performance of model outputs by comparing their win rates against the ground-truth response provided by the benchmark datasets.
-Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works [11, 29, 19], to perform the judgments.
+Considering the large amounts of test samples, we leverage GPT-4 [2], a widely utilized evaluation tool in previous works [11, 29, 19], to perform the judgments.
 Each target response, ground-truth response, query, and evaluated objectives are provided via prompt engineering.
-GPT-4 is required to compare and select the response with higher alignment to the specified objectives or indicate a tied performance.
-More details about the training process, model cards, dataset statistics, IMHI dataset annotation, and evaluation settings are presented in Appendix C.
+GPT-4 is required to compare and select the response with higher alignment on the specified objective or indicate a tied performance of the two responses.
+More details about the training process, model cards, dataset statistics, IMHI dataset annotation, and evaluation settings are presented in Appendix D.
+Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on 3 datasets over different policy models.
+The responses are simultaneously aligned on all trained objectives, then evaluated on each objective.
+"IF" denotes the "Instruction following" objective.
+"+" shows the advantage of aligned outputs over the unaligned outputs on win rates against the ground-truth responses.
 HH-RLHF UltraFeedback IMHI
@@
 Vicuna-33B +8.5% +2.59% +23.83% +8.0% +11.67% +6.33% +6.67% +8.34% +4.44% +6.12%
-GPT-3.5 +1.42% +7.5% +17.84% +5.0% +5.0% +3.66% +1.0% +9.67% +1.33% +9.33%
+GPT-3.5-Turbo +1.42% +7.5% +17.84% +5.0% +5.0% +3.66% +1.0% +9.67% +1.33% +9.33%
 Claude-3-Sonnet -3.83% +1.58% +13.17% +4.67% +2.67% +2.67% +3.0% +7.0% +2.33% +6.66%
@@
 Vicuna-33B +28.0% +17.09% +30.83% +30.0% +37.34% +32.33% +29.33% +11.11% +16.11% +8.34%
-GPT-3.5 +15.92% +21.5% +22.84% +29.99% +30.34% +28.0% +14.34% +18.67% +16.33% +14.22%
+GPT-3.5-Turbo +15.92% +21.5% +22.84% +29.99% +30.34% +28.0% +14.34% +18.67% +16.33% +14.22%
 Claude-3-Sonnet +19.17% +19.08% +26.17% +22.33% +21.0% +21.67% +19.0% +11.33% +19.33% +11.33%
-13B LLaMA2-Chat-7B +21.0% +30.5% +23.75% +51.83% +47.5% +45.33% +38.67% +28.33% +38.33% +50.56%
-LLaMA2-Chat-13B +22.75% +16.08% +12.25% +46.33% +48.67% +46.83% +41.17% +30.56% +37.22% +40.56%
-LLaMA2-Chat-70B +16.58% +11.92% +23.08% +44.33% +35.0% +45.5% +24.0% +31.67% +30.56% +36.12%
-Gemma-instruct-2B +27.5% +20.75% +24.33% +55.0% +44.67% +51.33% +36.83% +35.55% +63.33% +65.0%
-Gemma-instruct-7B +14.0% +27.75% +25.17% +42.0% +40.17% +35.17% +31.17% +34.45% +50.0% +49.44%
-Vicuna-7B +18.0% +12.83% +30.33% +41.5% +39.83% +44.33% +37.5% +24.44% +23.33% +21.11%
-Vicuna-13B +17.92% +29.0% +32.67% +47.33% +49.17% +47.0% +40.67% +28.33% +23.34% +18.9%
-Vicuna-33B +31.5% +17.09% +27.83% +50.5% +53.17% +45.83% +38.5% +23.89% +23.89% +14.45%
-GPT-3.5 +19.92% +25.0% +23.84% +40.33% +40.17% +36.83% +23.67% +26.67% +25.66% +33.62%
-Claude-3-Sonnet +27.17% +23.58% +27.17% +38.5% +39.5% +37.67% +29.83% +28.67% +20.0% +11.2%
-Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on 3 datasets over different policy models.
-The responses are simultaneously aligned on all objectives, then tested on each objective.
-"IF" denotes the "Instruction following" objective.
-"+" shows the advantage of aligned outputs over the unaligned outputs on win rates against the golden responses.
+13B LLaMA2-Chat-7B +24.0% +30.5% +23.75% +51.83% +47.5% +45.33% +38.67% +28.33% +38.33% +50.56%
+LLaMA2-Chat-13B +17.75% +16.58% +15.75% +46.33% +48.67% +46.83% +41.17% +30.56% +37.22% +40.56%
+LLaMA2-Chat-70B +16.58% +19.42% +26.58% +44.33% +35.0% +45.5% +24.0% +31.67% +30.56% +36.12%
+Gemma-instruct-2B +18.5% +17.25% +24.33% +55.0% +44.67% +51.33% +36.83% +35.55% +63.33% +65.0%
+Gemma-instruct-7B +17.5% +23.75% +30.17% +42.0% +40.17% +35.17% +31.17% +34.45% +50.0% +49.44%
+Vicuna-7B +19.0% +19.83% +26.33% +41.5% +39.83% +44.33% +37.5% +24.44% +23.33% +21.11%
+Vicuna-13B +18.92% +28.5% +32.67% +47.33% +49.17% +47.0% +40.67% +28.33% +23.34% +18.9%
+Vicuna-33B +31.5% +20.09% +27.83% +50.5% +53.17% +45.83% +38.5% +23.89% +23.89% +14.45%
+GPT-3.5-Turbo +18.42% +25.0% +29.34% +40.33% +40.17% +36.83% +23.67% +26.67% +25.66% +33.62%
+Claude-3-Sonnet +21.17% +20.58% +27.17% +38.5% +39.5% +37.67% +29.83% +28.67% +20.0% +11.2%
 4.2 Overall Performance
@@
 On the mental health analysis benchmark IMHI, MetaAligner models also show remarkable win rates on all objectives, proving their effectiveness in performing multi-objective alignment in domain knowledge-intense scenarios.
-We further evaluate MetaAligner on each IMHI sub-task and the results are shown in Appendix D.
-From the policy model scale perspective, MetaAligner provides successful alignments to open-sourced models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself.
-In the extreme case, MetaAligner-1.1B advances the win rates of LLaMA2-Chat-70B outputs, a policy model with 63$\times$ more parameters, by an average of 12.19% on HH-RLHF, 13.08% on UltraFedback, and 13.26% on IMHI.
+We further evaluate MetaAligner on each IMHI sub-task and the results are shown in Appendix E.
+From the policy model scale perspective, MetaAligner provides successful alignments to open-source models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself.
+In the extreme case, MetaAligner-1.1B advances the win rates of LLaMA2-Chat-70B outputs, a policy model with 63$\times$ more parameters, by an average of 12.19% on HH-RLHF, 13.08% on UltraFeedback, and 13.26% on IMHI.
 These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow.
-MetaAligner also significantly improves performance on API-based models: GPT-3.5 and Claude-3-Sonnet.
-These results prove its potential for application in close-sourced scenarios and effective multi-objective alignment of state-of-the-art policy models.
-Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase.
-These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner.
+MetaAligner also significantly improves performance on close-source LLMs: GPT-3.5-Turbo and Claude-3-Sonnet.
+These results prove its potential for application in close-source scenarios and effective multi-objective alignment of state-of-the-art policy models.
+Within most policy model families, we observe a decreasing trend in win-rate advantage as their sizes increase.
+These decreases indicate a struggle aligning powerful large-scale policy models with small MetaAligner models.
 Fortunately, MetaAligner’s capabilities also show scalability.
@@
 For example, on UltraFeedback, MetaAligner-7B outperforms MetaAligner-1.1B on all 10 policy models, and MetaAligner-13B further surpasses MetaAligner-7B by an average of 12.58%.
-These observations motivate further explorations in model scale-performance balance for MetaAligner.
-4.3 Evaluations of Objective-Wise Alignment
-Figure 2: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
-The results are the performance of MetaAligner-7B on LLaMA2-Chat-70B outputs from the UltraFeedback test set.
-We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives.
-We utilize MetaAligner to perform 6 levels of alignments: unaligned, aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and full alignment.
-We leverage GPT-4 to score the responses ranging from 0 to 10.
-Experimental details and more results are shown in Appendix E.
-The results are shown in Figure 2.
-We have the following observations:
-Objective-wise alignment improves performance on the target objective and benefits the performance on other objectives.
-For example, Aligning on "Instruction following" achieves the best GPT4 score distribution on the "Instruction following" evaluation results.
-It also significantly increases GPT4 scores on "Helpful", "Honest", and "Truthful" over the unaligned responses.
-This tendency holds with other policy models and alignment objectives.
-These results further prove the complex interplay within multi-objective alignment, where positive correlations and contradictions [11] co-exist.
-Full alignment on all objectives provides balanced performance.
-According to the results, full alignment displays competitive performance on all 4 objectives.
-Generally, it outperforms unaligned outputs and aligned outputs from other objectives, even comparable to those from the same objective, such as in "Honest".
-The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, training the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
-Figure 3: Zero-shot alignment on 6 unseen objectives based on LLaMA2-Chat-70B outputs of UltraFeedback.
+These observations motivate further explorations in model size-performance balance for MetaAligner.
+Table 3: Comparisons of win rates between alignment methods.
+"GPU Hours" records the summed GPU running time on all datasets.
+"-Equal Pref." and "-Warm Up" denote the removal of the "equal-preference alignment" and "warming up" stages.
+HH-RLHF UltraFeedback
+Policy Model Algorithm GPU Hours Harmless Helpful Humour Avg. IF Honest Truthful Helpful Avg.
+LLaMA2-Chat-7B MORLHF 1892.3 62.83% 51.2% 77.5% 63.84% 32.18% 33.7% 26.1% 33.7% 31.42%
+MODPO 405.9 65.0% 64.0% 78.0% 69.0% 30.82% 43.4% 37.19% 25.0% 34.1%
+SFT 247.34 66.5% 75.0% 76.5% 72.67% 27.0% 36.5% 26.0% 36.5% 31.5%
+Aligner-7B 236.8 72.0% 81.9% 70.12% 74.67% 52.38% 44.23% 37.19% 39.1% 43.23%
+MetaAligner-1.1B 120.48 62.5% 75.0% 77.0% 71.5% 27.67% 27.0% 33.0% 25.33% 28.25%
+MetaAligner-7B 242.68 77.5% 82.0% 83.0% 80.83% 51.33% 48.0% 55.67% 44.33% 49.83%
+-Equal Pref. 73.82% 80.7% 77.39% 77.3% 46.8% 43.6% 53.17% 41.7% 46.32%
+-Warm Up 77.1% 80.32% 82.63% 80.02% 49.96% 47.4% 55.73% 44.18% 49.32%
+MetaAligner-13B 403.44 76.5% 85.5% 86.0% 82.67% 68.5% 59.5% 64.0% 55.0% 61.75%
+LLaMA2-Chat-70B Self-Refinement 70.48% 82.8% 68.91% 74.06% 49.95% 62.91% 60.77% 57.6% 55.05%
+MetaAligner-7B 242.68 85.16% 89.42% 88.08% 87.55% 67.05% 63.72% 70.1% 54.7% 63.89%
+4.3 MetaAligner vs. Baseline Methods
+We compare the performance of MetaAligner with MORLHF, MODPO, SFT-based methods, and Aligner.
+We implement the linear scalarization method for MORLHF, the CDPO [11] realization of MODPO, and RiC [36] realization of the SFT-based method.
+As Aligner is not suitable for multi-objective alignment, we train Aligner-7B on "Helpful" annotations for HH-RLHF and "IF" for UltraFeedback.
+We compare these methods on the LLaMA2-Chat-7B policy model.
+We further include a self-refinement method which prompts the policy model itself to refine its own outputs.
+We compare self-refinement on the LLaMA2-Chat-70B policy model as it requires strong in-context learning ability from the policy model.
+The results are presented in Table 3.
+Appendix G presents details about the baseline model implementations and GPU hours calculations.
+According to the results, MetaAligner-13B significantly outperforms all other methods with an average of 82.67% win rate on HH-RLHF and 61.75% on UltraFeedback, showing the general advantage of the conditional weak-to-strong correction paradigm.
+As the base model size reduces, MetaAligner shows decreased but still competitive performance compared to other baseline models, but achieved with less memory consumption and GPU training hours.
+Impressively, the MetaAligner-1.1B model achieves comparable average performance to MORLHF, MODPO, and SFT-based methods on both datasets, but costs only 6.37%-48.71% of their GPU training hours, with a 6.36$\times$ smaller size than the LLaMA2-Chat-7B policy model.
+These facts indicate the high efficiency of MetaAligner algorithms and a prospect for application in low-resource scenarios.
+Compared to previous methods, MetaAligner models can also achieve balanced and stable performances in objective-wise evaluations, including contradictory objectives such as "Harmless" and "Helpful", without requiring explicit hyper-parameter tuning for achieving Pareto optimal solutions [11, 36, 20].
+In other methods, inappropriate heuristic preference weight selection can lead to serious performance degradation in certain objectives.
+For example, with a uniform distribution of preference weights, the performance of MORLHF on "Helpful" falls to 51.2%, a huge gap to other methods.
+Though Aligner-7B is comparable to MetaAligner-7B on its aligned objectives "Helpful" and "IF", it significantly underperforms MetaAligner in other objectives.
+These results prove the effectiveness of MetaAligner in simultaneously aligning multiple objectives.
+MetaAligner-7B also outperforms self-refinement with the LLaMA2-Chat-70B policy model on 6 of 7 objectives with only 1/10 in inference cost, showing the necessity of training specific modules for multi-objective alignment.
+Ablation studies on MetaAligner-7B show that both warming up and equal-preference alignment stages make considerable contributions to model performance, with the removal of equal-preference alignment leads to a substantial decrease of 3.53% on HH-RLHF and 3.51% on UltraFeedback in average win rates.
+Figure 3: Zero-shot alignment on 6 unseen objectives.
 In the x-axis, "Aligned Obj." denotes the 4 supervised objectives ($\diamond$" markers), and "+" denotes further addition of an unseen objective ($\circ$" markers).
-$\star$" denotes the unaligned win rates for the unseen objectives, "-." lines identify win rate fluctuations before alignment, and solid lines identify fluctuations after alignment.
-4.4 Generalizable Alignment
-In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 6 less relevant objectives that are unseen during its training phases: Specific, Factual, Readable, Fair, Repeat, and Length [10].
-More details about these objectives are in Appendix B.
+$\star$" denotes the win rates for the unseen objectives before all zero-shot alignments, "-." lines identify win rate fluctuations before alignment, and solid lines identify fluctuations after alignment.
+4.4 Generalizable Alignment to Unseen Objectives
+In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with six unseen objectives: "Specific", "Factual", "Readable", "Fair", "Repeat", and "Length" [10].
+More details about these objectives are in Appendix C.
 We randomly select 2,700 queries from the UltraFeedback dataset and re-align the LLaMA2-Chat-70B outputs with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one, with 10 aligned objectives in total.
 Their win rates on each objective over the golden responses are presented in Figure 3.
-According to the results, we have the following conclusions:
+We have the following conclusions:
 MetaAligner performs effective zero-shot alignment for unseen objectives.
@@
 However, the win rates on supervised objectives ("Instruction following", "Helpful", "Honest", and "Truthful") generally surpass unseen objectives, showing that supervised learning remains more effective in multi-objective preference alignment compared to in-context learning.
-Performance on aligned objectives is maintained with more unseen alignment objectives.
+Performance on aligned objectives is maintained with additional unseen alignment objectives.
 As each objective is aligned, its win rate surges, stabilizing as long as it is included.
@@
 These results prove MetaAligner to perform overall reliable alignment with the expansion of objectives.
-However, aligning on new objectives can affect other objectives.
-For example, aligning on "Fair" with MetaAligner-(7B, 13B) benefits its win rates, but harms performance on objectives such as "Readable" and "Factual".
-The decreases are due to the "alignment tax" [11], where enhancements in one objective can diminish performance in other objectives due to their controversial nature.
+However, enhancements in one objective can affect performance in certain objectives due to their controversial nature, which is known as the "alignment tax" [11].
+For example, aligning on "Fair" (+Fair) with MetaAligner-(7B, 13B) benefits its win rates, but harms performance on objectives such as "Readable" and "Factual" compared to when "Fair" is unaligned.
 MetaAligner’s generalizability shows scalability.
-Performance on the 6 unseen objectives increases with the scale-up of MetaAligner model size.
-MetaAligner-1.1B provides limited alignment on most unseen objectives, but MetaAligner-7B extends the win rates to an average of 48.5%, and MetaAligner-13B further reaches 61.25%.
+Performance on the six unseen objectives increases with the scale-up of MetaAligner model size.
+MetaAligner-1.1B provides limited improvement on most unseen objectives, but MetaAligner-7B extends the win rates to an average of 48.5%, and MetaAligner-13B further reaches 61.25%.
 MetaAligner-13B also more effectively aligns objectives such as "Length", where smaller models perform badly.
 This scalability is attributed to larger foundation models’ growing in-context learning ability, which enables accurate interpretations of the objective descriptions and instructions.
-These observations motivate further explorations into the correlation between generalizable alignment and model scales in future work.
-HH-RLHF UltraFeedback
-Algorithm GPU Hours Harmless Helpful Humour Avg. IF Honest Truthful Helpful Avg.
-MORLHF 1892.3 62.83% 51.2% 77.5% 63.84% 32.18% 33.7% 26.1% 33.7% 31.42%
-MODPO 405.9 65.0% 64.0% 78.0% 69.0% 30.82% 43.4% 37.19% 25.0% 34.1%
-SFT 247.34 66.5% 75.0% 76.5% 72.67% 27.0% 36.5% 26.0% 36.5% 31.5%
-MetaAligner-1.1B 120.48 62.5% 75.0% 77.0% 71.5% 27.67% 27.0% 33.0% 25.33% 28.25%
-MetaAligner-7B 242.68 77.5% 82.0% 83.0% 80.83% 51.33% 48.0% 55.67% 44.33% 49.83%
-MetaAligner-13B 403.44 73.5% 85.5% 86.0% 81.67% 68.5% 59.5% 64.0% 55.0% 61.75%
-Table 3: Comparisons of win rates between alignment methods based on the LLaMA2-Chat-7B policy model. "GPU hours" records the summed GPU running time on all datasets.
-4.5 MetaAligner vs. MORLHF/MODPO/SFT-based Method
-We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on the HH-RLHF and UltraFeedback datasets, where explicit rewards are available.
-We implement the linear scalarization method for MORLHF, the CDPO [11] realization of MODPO, and RiC [35] realization of the SFT-based method.
-We select LLaMA2-Chat-7B as the policy model, and the results are presented in Table 3.
-Appendix F presents details about the baseline model implementations and GPU hours calculations.
-According to the results, MetaAligner-13B significantly outperforms all other methods with an average of 81.67% win rate on HH-RLHF and 61.75% on UltraFeedback, showing the general advantage of the conditional weak-to-strong correction paradigm.
-As the base model size reduces, MetaAligner shows decreased but still competitive performance compared to other baseline models, but achieved with less memory consumption and GPU training hours.
-Impressively, the MetaAligner-1.1B model achieves comparable average performance to MORLHF, MODPO, and SFT-based methods on both datasets, but with 2.05$\times$-15.67$\times$ less GPU training hours and 6.36$\times$ smaller size than the LLaMA2-Chat-7B policy model.
-These facts indicate the high efficiency of MetaAligner algorithms and a prospect for application in low-resource scenarios.
-Compared to previous methods, MetaAligner models can also achieve balanced performances in objective-wise evaluations, including contradictory objectives such as "Harmless" and "Helpful", without requiring explicit hyper-parameter tuning for achieving Pareto optimal solutions [11, 35, 20].
-In other methods, inappropriate heuristic preference weight selection can lead to serious performance degradation in certain objectives.
-For example, with a uniform distribution of preference weights, the performance of MORLHF on "Helpful" falls to 51.2%, a huge gap to other methods.
+These observations motivate further explorations into the correlation between generalizable alignment and base model scales in future work.
+4.5 Evaluations of Objective-Wise Alignment
+Figure 4: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
+The results are the performance of MetaAligner-7B on LLaMA2-Chat-70B outputs from the UltraFeedback test set.
+We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives.
+We utilize MetaAligner to perform six levels of alignments: unaligned, aligning on each objective ("Instruction following", "Helpful", "Honest", and "Truthful"), and full alignment.
+We leverage GPT-4 to score the responses ranging from 0 to 10.
+The results are shown in Figure 4.
+Experimental details and more results are shown in Appendix F.
+We have the following observations:
+Objective-wise alignment improves performance on the primary target and boosts the performance on other objectives.
+For example, Aligning on "Instruction following" achieves the best GPT4 score distribution on the "Instruction following" evaluation results.
+It also significantly increases GPT4 scores on "Helpful", "Honest", and "Truthful" over the unaligned responses.
+This tendency holds with other policy models and alignment objectives.
+These results further prove the complex interplay among objectives, where correlations and contradictions [11] co-exist.
+Full alignment on all objectives provides balanced performance.
+According to the results, full alignment displays competitive performance on all 4 objectives.
+Generally, it outperforms unaligned outputs and aligned outputs from other objectives, even comparable to those from the same objective, such as in "Honest".
+The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, training the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
 5 Related Work
-5.1 Large Language Models
-Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language [2, 31, 21].
-Notable examples include commercially available LLMs like ChatGPT, GPT-4 [2], and Claude-3 [3], as well as leading-edge open-source models like LLaMA2 [31] and Mistral [15].
-Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information [37, 32], flattery, and deception, raising concerns about their potential negative impacts on society and associated risks [5].
-To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning [14].
-5.2 Human Preference Alignment
-Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty.
-These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) [23].
-Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences.
-This limitation has motivated the exploration of multi-objective RLHF (MORLHF) [27, 20, 25], by linear scalarizations of multiple rewards [27, 20] and interpolations of LLM weights trained from diverse reward models [25].
-However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) [38] without the RL paradigm.
-Multi-objective direct preference optimization (MODPO)[38] extended DPO algorithm [24] to combine multiple objectives with specific weightings.
-Another paradigm is the SFT-based methods [35, 11], which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO.
-These methods also facilitated objective-wise controllable generation during inference.
+This paper focuses on advancing multi-objective alignment of language models with human values, which is mainly related to two research areas: (1) Large Language Models, including the latest development in close-source AI agents [2, 22, 3] and open-source foundation models [31, 30, 7].
+(2) Alignment of Language Models, including RLHF [41, 23, 28] and its enhanced variants [24, 37, 13].
+Multi-objective alignment methods include MORLHF [27, 20, 25], MODPO [40, 11], and SFT-based methods [36, 11].
+A detailed review of related work is in Appendix B.
 6 Discussions
@@
 This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment.
-Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives.
-Thorough investigations on various LLM policy models proved MetaAligner’s overall effectiveness in multi-objective alignment and objective-wise alignment.
+It follows a three-stage training paradigm: (1) dynamic objectives reformulation; (2) conditional weak-to-strong correction; (3) generalizable inference for multi-objective alignment.
+MetaAligner can perform plug-and-play inference and zero-shot alignment to unseen objectives.
+Thorough investigations on various policy models proved MetaAligner’s overall effectiveness in multi-objective and objective-wise alignment.
 Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align multiple objectives.
-Limitation and Future Work.
+Limitations and Future Work.
 Firstly, stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference [13], which affects model deployment, especially for scenarios such as local deployment on mobile devices.
@@
 We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 10 objectives).
-This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
+It will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
+Acknowledgements
+This work is supported by the computational shared facility and President’s Doctoral Scholar award, The University of Manchester.
+This work is supported by the project JPNP20006 from New Energy and Industrial Technology Development Organization (NEDO), and Artificial Intelligence Research Center, National Institute of Advanced Industrial Science and Technology, Japan.
 Appendix A Ethics and Impacts
@@
 In Sec.
-4.5, we implement the reward assignment scripts for HH-RLHF and MORLHF based on the released codes of Yang et al.
+4.3, we implement the reward assignment scripts for HH-RLHF and MORLHF based on the released codes of Yang et al.
 [36], which is available at Github.
@@
 By using or accessing the information in this paper, the users agree to indemnify, defend, and hold harmless the authors, contributors, and any affiliated organizations or persons from any and all claims or damages.
-Appendix B Text Descriptions and Prompting Templates
+Appendix B Related Work
+B.1 Large Language Models
+Large language models (LLMs) have reached approaching-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language [2, 31, 21].
+Notable examples include commercially available LLMs like ChatGPT [22], GPT-4 [2], and Claude-3 [3].
+Due to the high inference cost of these close-source models, research trend in open-source foundation models surges, leading to cutting-edge open-source models like LLaMA2 [31], Gemma [30], and Vicuna [7].
+Open-source models, though underperforming state-of-the-art commercial models in instruction following and reasoning capabilities, provide fully accessible model parameters to facilitate efficient inference and customized parameter fine-tuning.
+Despite the advancements of LLMs, recent studies found that they can exhibit problematic behaviors, including the generation of inaccurate information [39, 33], flattery, and deception, raising concerns about their potential negative impacts and associated risks on society [5].
+To address these issues, considerable research has been dedicated to refining LLMs’ outputs to better align with human values and preferences [14].
+B.2 LLM Alignment on Human Values
+Many studies have delved into enhancing the responses of LLMs in core characteristics of human values like "Helpful", "Harmless", and "Honest".
+Early efforts are largely centered on Reinforcement Learning from Human Feedback (RLHF) [41, 23, 28], where the alignment of human values is manifested by maximizing a scalar value obtained from the reward model with a KL-regularization, using RL-algorithms such as PPO [26].
+However, PPO faces challenges including inefficiency and instability, driving development in simplified algorithms such as DPO [24], rank-based learning [37], and weak-to-strong correction [13].
+Nonetheless, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, where a scalar reward model proves inadequate for aligning LLMs with varied human preferences.
+This limitation motivates the exploration of more complex alignment objectives, including fine-grained human feedbacks [32, 17, 10] via reward value breakdown or compositions, and multi-objective preference alignment.
+Some works explored multi-objective RLHF (MORLHF) [27, 20, 25], by linear scalarizations of multiple rewards [27, 20] or interpolations of LLM weights trained from diverse reward models [25].
+However, diverse reward models can increase the computational cost, and the PPO training paradigm still leads to training challenges due to its unstable nature.
+Recent studies further explore the multi-objective direct preference optimization (MODPO) [40, 11] without the RL paradigm.
+MODPO extends the DPO algorithm to combine multiple objectives with specific weightings [40] or controllable preference values [11].
+Other methods further simplify the paradigm with SFT-based methods [36, 11], which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO.
+These methods also facilitated objective-wise controllable generation during inference.
+Compared with previous methods, MetaAligner performs conditional weak-to-strong correction based on the SFT paradigm, which offers reward-free, policy model-agnostic, and flexible multi-objective preference alignment.
+The model also effectively aligns unseen objectives, mark- ing the first step towards generalizable multi-objective preference alignment.
+Appendix C Details of Dynamic Objectives Reformulation
+C.1 Objective Descriptions
 The text descriptions for all tested objectives in this paper are included in Table 4.
@@
 The aligned objectives are obtained from the annotations of each alignment dataset, and the unaligned objectives are selected from the definitions of previous works [10].
+C.2 Prompting Templates
 On building the dynamic multi-objective dataset, we carefully define prompting templates to trigger the conditional weak-to-strong correction ability of MetaAligner.
@@
 Consider the query, then edit the answer to make it equal considering these aspects: $\mathcal{O}$} | Query: {q} | Answer: {y} | Edit:
-Objectives $r$ Text Description $r\right>$
+Table 4: Text descriptions for all tested objectives.
+Objectives Text Description $d\right>$
 Aligned Objectives
-Harmless The response should avoid content that is offensive, discriminatory, or harmful.
-Helpful The response should provide useful resources and suggestions to the user.
-Humor The response should be cheerful and amusing.
-Instruction following The response should carefully follow the instructions of the query.
-Honest The response should not tell lies
-Truthful The response should actively make known all the full truth of a matter
-Correct The explanations should make correct predictions.
-Informative The response should express clear logic and provide consistent evidence.
-Professional The response should provide evidence with high quality and reliability.
+Harmless Harmlessness: The response should avoid content that is offensive, discriminatory, or harmful.
+Helpful Helpfulness: The response should provide useful resources and suggestions to the user.
+Humor Humor: The response should be cheerful and amusing.
+Instruction following Instruction following: The response should carefully follow the instructions of the query.
+Honest Honesty: The response should not tell lies
+Truthful Truthfulness: The response should actively make known all the full truth of a matter
+Correct Correctness: The explanations should make correct predictions.
+Informative Informativeness: The response should express clear logic and provide consistent evidence.
+Professional Professionalism: The response should provide evidence with high quality and reliability.
 Unaligned Objectives
-Specific The response should refer to facts and details and avoid vague arguments.
-Factual The response should be factually correct and avoid hallucinated statements.
-Readable The response should be easy to read and understand, not too technical for laymen.
-Fair The response should avoid biased or one-sided arguments and consider different points of view.
-Repeat The response should avoid repetitive statements of one point.
-Length The response should be concise and avoid redundant content.
-Table 4: Text descriptions for all tested objectives.
-Appendix C Experimental Details
-C.1 Model Training
+Specific Specificity: The response should refer to facts and details and avoid vague arguments.
+Factual Factuality: The response should be factually correct and avoid hallucinated statements.
+Readable Readability: The response should be easy to read and understand, not too technical for laymen.
+Fair Fairness: The response should avoid biased or one-sided arguments and consider different points of view.
+Repeat Repetition: The response should avoid repetitive statements of one point.
+Length Length: The response should be concise and avoid redundant content.
+C.3 An Example of Dynamic Objectives Reformulation
+We provide an example of the dynamic objectives reformulation algorithm 1 to better illustrate its working mechanism.
+We use one instance in the HH-RLHF dataset as the example and focus on one running of the loop in line 2.
+Specifically, we have
+$\mathbf{q}$: Human: What are some cool countries to visit in Asia?
+Assistant: Hm, it’s difficult to pick just one.
+Thailand, Japan, Vietnam, Indonesia, and many others have unique history and culture, and excellent beaches and good food.
+Also many Asian countries are rapidly growing and developing, with good infrastructure, and many people want to see what’s happening there, from my observation.
+There are also many cities around the world that would have more to see, like Bangkok and Shanghai.
+There are also cities that I’ve visited in Asia, like Xi’an, which I would really recommend.
+Human: Okay, which one do you like?
+Assistant: $\mathbf{y_{1}}$: If I was a tourist in Asia, I’d love to visit Kolkata.
+Maybe Xi’an, because it’s just a cool place to visit.
+$\mathbf{y_{2}}$: I like Thailand and Japan, because they have a strong culture that is still being preserved and a unique history.
+We have the following reward score assignments from the reward models:
+Response Helpful Harmless Humour
+$y_{1}$ 0.12 2.15 0.43
+$y_{2}$ 0.37 2.17 0.29
+We can map the above reward values to preferences on each objective: $P=[\prec,\prec,\succ]$.
+Based on this preference vector, we obtain the text description sets after the processing on line 5-12, which are as follows:
+[$\mathcal{O}_{\succ}$] Humor: The response should be cheerful and amusing;
+[$\mathcal{O}_{\prec}$] Harmless: The response should avoid content that is offensive, discriminatory, or harmful; Helpful: The response should provide useful resources and suggestions to the user;
+[$\mathcal{O}_{\equiv}$] $\varnothing$
+Based on the above information, we can build two pairs of weak-to-strong training samples, by fitting in the templates of HH-RLHF dataset provided in Sec. C.2:
+QUERY 1: You are an assistant to human. You will be provided with a context and an answer. Consider the context, then edit the answer to improve it considering these aspects:
+Harmlessness: The response should avoid content that is offensive, discriminatory, or harmful;
+Helpfulness: The response should provide useful resources and suggestions to the user | Context: Human: What are some cool countries to visit in Asia?
+Assistant: Hm, it’s difficult to pick just one.
+Thailand, Japan, Vietnam, Indonesia, and many others have unique history and culture, and excellent beaches and good food.
+Also many Asian countries are rapidly growing and developing, with good infrastructure, and many people want to see what’s happening there, from my observation.
+There are also many cities around the world that would have more to see, like Bangkok and Shanghai.
+There are also cities that I’ve visited in Asia, like Xi’an, which I would really recommend.
+Human: Okay, which one do you like?
+Assistant: | Answer: If I was a tourist in Asia, I’d love to visit Kolkata.
+Maybe Xi’an, because it’s just a cool place to visit.
+| Edit:
+CORRECTION 1:
+I like Thailand and Japan, because they have a strong culture that is still being preserved and a unique history.
+QUERY 2: You are an assistant to human. You will be provided with a context and an answer. Consider the context, then edit the answer to improve it considering these aspects:
+Humor: The response should be cheerful and amusing |
+Context: Human: What are some cool countries to visit in Asia?
+Assistant: Hm, it’s difficult to pick just one.
+Thailand, Japan, Vietnam, Indonesia, and many others have unique history and culture, and excellent beaches and good food.
+Also many Asian countries are rapidly growing and developing, with good infrastructure, and many people want to see what’s happening there, from my observation.
+There are also many cities around the world that would have more to see, like Bangkok and Shanghai.
+There are also cities that I’ve visited in Asia, like Xi’an, which I would really recommend.
+Human: Okay, which one do you like?
+Assistant: | Answer: I like Thailand and Japan, because they have a strong culture that is still being preserved and a unique history.
+| Edit:
+CORRECTION 2:
+If I was a tourist in Asia, I’d love to visit Kolkata. Maybe Xi’an, because it’s just a cool place to visit.
+Since $\mathcal{O}_{\equiv}$ is empty, this data sample does not contribute to the equal subset $\mathcal{D}_{e}$.
+The two created pairs are incorporated into the dynamic multi-objective dataset as two instances.
+Appendix D Experimental Details
+D.1 Model Training
 Details about the training process of MetaAligner and the building process of the 3 datasets are presented in Table 5.
@@
 With the above information and the target objective description $r_{t}\right>$, we obtain the win rates using Algorithm 2.
-C.2 Model Cards
+D.2 Model Cards
 TinyLLaMA-1.1B [38].
@@
 MentaLLaMA can perform mental health analysis on social media data and generate high-quality explanations for its predictions.
-On evaluating sub-task performance on IMHI Benchmark (Appendix D), we introduce MentaLLaMA-(7B, 13B, 33B) models as domain-specific policy models to evaluate the alignment performances.
-GPT-3.5 [22].
-GPT-3.5 is an advanced, close-sourced chat-based language model developed by OpenAI.
+On evaluating sub-task performance on IMHI Benchmark (Appendix E), we introduce MentaLLaMA-(7B, 13B, 33B) models as domain-specific policy models to evaluate the alignment performances.
+GPT-3.5-Turbo [22].
+GPT-3.5-Turbo is an advanced, close-source chat-based language model developed by OpenAI.
 It is a sibling model to InstructGPT, which is trained to follow instructions in a prompt and provide a detailed response.
@@
 IMHI-MetaAligner focuses on the interpretable mental health analysis domain and is trained to align responses of an AI psychologist on analyzing mental health conditions based on social media posts.
-C.3 IMHI Annotation
+Figure 5: Heatmap of the objective distributions on the dynamic multi-objective dataset built from IMHI.
+"Pro." and "Info." denote the "Professional" and "Informative" objectives.
+Similarly, the dataset also shows a balanced overall distribution and objective-wise distribution.
+D.3 IMHI Annotation
 We select 1,200 queries from the IMHI benchmark covering 9 mental health analysis tasks.
@@
 26:$\omega=\omega_{aligned}-\omega_{origin}$
+Table 5: Details for MetaAligner training and datasets. ‘preference source’ denotes how the preference annotations were obtained.
 Training Information
@@
 Dataset Name HH-RLHF
-Licence MIT
+License MIT
 Train/Val/Test ($\mathcal{D}_{p}$) 262,719/15,000/15,000
@@
 Dataset Name UltraFeedback
-Licence MIT
+License MIT
 Train/Val/Test ($\mathcal{D}_{p}$) 252,934/15,000/15,000
@@
 Dataset Name IMHI
-Licence MIT
+License MIT
 Train/Val/Test ($\mathcal{D}_{p}$) 5,304/1,051/2,400
@@
 Vicuna-(7B, 13B, 33B) https://huggingface.co/lmsys
-GPT-3.5 https://openai.com/blog/chatgpt
+GPT-3.5-Turbo https://openai.com/blog/chatgpt
 Claude-3-Sonnet https://www.anthropic.com/news/claude-3-family
-Table 5: Details for MetaAligner training and datasets. ‘preference source’ denotes how the preference annotations were obtained.
-Appendix D Sub-task Performance on IMHI Benchmark
+Appendix E Sub-task Performance on IMHI Benchmark
+Table 6: Details about the 9 sub-tasks in the IMHI dataset. "Annotation" denotes the reliability of the annotations in the raw data.
 Data Task Source Annotation Labels/Aspects
@@
 IRF interpersonal risk factors detection Reddit human annotation Thwarted Belongingness, Perceived Burdensomeness
-Table 6: Details about the 9 sub-tasks in the IMHI dataset. "Annotation" denotes the reliability of the annotations in the raw data.
+Table 7: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models.
+The GPT-4 judge considers 3 objectives: "Correct", "Informative", and "Professional".
+The figures show the advantage of aligned outputs over the policy model outputs on win rate.
+Best values for each MetaAligner model are highlighted in bold.
 MetaAligner Policy Model CAMS DR Dreaddit IRF loneliness MultiWD SAD SWMH T-SID Overall
@@
 Vicuna-33B -6.0% +54.0% +36.0% +79.0% +38.0% +29.0% +24.0% +21.0% +24.0% +33.22%
-Table 7: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models.
-The GPT-4 judge considers 3 objectives: "Correct", "Informative", and "Professional".
-The figures show the advantage of aligned outputs over the policy model outputs on win rate.
-Best values for each MetaAligner model are highlighted in bold.
 We stack MetaAligner on different policy models to perform alignment on all 3 objectives: "Correct", "Informative", and "Professional".
@@
 These results show that MetaAligner can make reasonable corrections on weak responses while maintaining their expertise from domain-specific policy models.
-Appendix E Objective-wise Alignment
-E.1 Experimental Settings
+Appendix F Objective-wise Alignment
+F.1 Experimental Settings
 We randomly sample 1,200 queries from the UltraFeedback test set and probe the target policy models to provide responses to all queries, which are regarded as unaligned responses.
@@
 where $q$ and $a$ denotes the target query and response, and $r\in${Instruction following, Helpful, Honest, Truthful} is the target objective for evaluation.
-E.2 Experimental Results
+F.2 Experimental Results
 More experimental results are presented in Figure 6.
-The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5 outputs from the UltraFeedback test set.
-According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5.
+The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5-Turbo outputs from the UltraFeedback test set.
+According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5-Turbo.
 Unlike full alignment which shows weaker performance as the capability of the policy model increases, objective-wise alignment provides stable improvement for different policy models.
@@
 Figure 6: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
-Appendix F Implementation Details of Baseline Models
-F.1 SFT-based Methods
+Appendix G Implementation Details of Baseline Models
+G.1 SFT-based Methods
 Existing SFT-based methods [11, 36] for multi-objective preference alignment explicitly incorporate reward values into the query via prompt engineering, which includes a text marker for each alignment objective and their corresponding value numbers in the current response.
@@
 In calculating the GPU hours, we sum the training hours for fine-tuning the LLaMA-Chat-7B policy model for HH-RLHF and UltraFeedback datasets.
-F.2 MODPO
+G.2 MODPO
 We implement MODPO by implementing the Controllable Direct Preference Optimization (CDPO) [11] algorithm based on its paper introduction.
@@
 where $\omega_{i}$ represents the weight of the $i$-th objective and $g_{i}$ is calculated as follows:
-$\small g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (10)
+$g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (8)
 where $\lambda_{i}$ represents the weight of the controlled objective and $p_{j}$ is a pre-defined preference value for the $j$-th objective.
@@
 We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for fair comparisons.
-F.3 MORLHF
+G.3 MORLHF
 We use the linear scalarization [27, 20] realization of MORLHF with the KL-divergence regularization. Specifically, the model is trained via the following objective function:
@@
 Finally, the PPO training hours for HH-RLHF and UltraFeedback are included in the GPU hours.
+G.4 Self-Refinement
+We include a prompt engineering-based self-refinement method as a baseline method to further demonstrate the effectiveness of MetaAligner.
+Specifically, this approach involves obtaining an initial response from the policy model and then prompting the same model to further refine its own output.
+For the second stage, we utilize the same prompting strategies as MetaAligner, which is detailed in Appendix C.2.
+However, this method often requires aligner models with strong in-context learning capabilities, leading to high inference costs due to larger model sizes or expensive commercial models.
+Therefore, we select LLaMA2-Chat-70B, a strong policy model as the target policy model for self-refinement.
+As self-refinement does not involve any training procedures, we do not report its GPU hours as other methods.